### v2

# Counting Small Induced Subgraphs:

Abstract
For a fixed graph property $\Phi$ and integer $k\geq 1$, the problem #$\#\textnormal{{IndSub}}(\Phi,k)$ asks to count the induced $k$-vertex subgraphs satisfying $\Phi$ in an input graph $G$. If $\Phi$ is trivial on $k$-vertex graphs (i.e., if $\Phi$ contains either all or no $k$-vertex graphs), this problem is trivial. Otherwise we prove, among other results:
  • If $\Phi$ is edge-monotone (i.e., closed under deleting edges), then #$\#\textnormal{{IndSub}}(\Phi,k)$ cannot be solved in time $n^{o(k)}$ assuming ETH. This strengthens a result by Döring, Marx and Wellnitz [STOC 2024] that only ruled out an exponent of $o(\sqrt{\log k}/\log\log k)$. Our results also hold when counting modulo fixed primes.
  • If there is some fixed $\varepsilon>0$ such that at most $(2-\varepsilon)^{\binom{k}{2}}$ graphs on $k$ vertices satisfy $\Phi$, then #$\#\textnormal{{IndSub}}(\Phi,k)$ cannot be solved in time $n^{o(k/\sqrt{\log k})}$ assuming ETH. Our results hold even when each of the graphs in $\Phi$ may come with an arbitrary individual weight. This generalizes previous results for hereditary properties by Focke and Roth [SIAM J. Comput. 2024] up to a $\sqrt{\log k}$ factor in the exponent of the lower bound.
  • If $\Phi$ only depends on the number of edges, then #$\#\textnormal{{IndSub}}(\Phi,k)$ cannot be solved in time $n^{o(k)}$ assuming ETH. This improves on a lower bound by Roth, Schmitt and Wellnitz [FOCS 2020] that only ruled out an exponent of $o(k/\sqrt{\log k})$.
In all cases, we also obtain #W[1]-hardness if $k$ is part of the input and the problem is parameterized by $k$. We also obtain lower bounds on the Weisfeiler-Leman dimension.
Most importantly, our results follow from relatively straightforward Fourier analysis, as opposed to the nontrivial techniques from combinatorics, group theory, and simplicial topology used in previous papers. Our paper subsumes most of the known #W[1]-hardness results known in the area, often with tighter lower bounds under ETH.
{textblock}
5(8.5, 8.75) [Uncaptioned image]
1 Introduction
Counting small patterns in graphs is of fundamental importance in computer science. In network analysis and bioinformatics, the counts of small patterns give insights into properties of the network, e.g., when a pattern count deviates significantly from the expected value in a random graph [34]. Counting patterns is also interesting from a theoretical perspective: For example, counting perfect matchings in bipartite graphs is precisely the problem of evaluating permanents of zero-one matrices; permanents play a central role in counting complexity and algebraic complexity theory [47, 1].
Counting Small Induced Subgraphs
While counting perfect matchings amounts to counting subgraph patterns $H$ whose size is comparable to the host graph $G$, many practical applications ask us to count patterns $H$ that are much smaller than $G$. Such problems can be formalized and analyzed in the following ways:
  1. 1.
    For a fixed pattern $H$, consider the function #$\#\mathrm{IndSub}(H\to\,\star\,)$: Given an input graph $G$, this function outputs the number #$\#\mathrm{IndSub}(H\to G)$ of induced $H$-copies in $G$. Examples are the numbers of triangles or $K_{4}$-copies in a graph $G$. While each such function can be computed in polynomial time, we wish to use hardness assumptions such as the Exponential-Time Hypothesis (ETH) [25] or other complexity assumptions to bound the exponent of this polynomial running time, in dependence on $H$.
  2. 2.
    We may also view both $H$ and $G$ as input when computing #$\#\mathrm{IndSub}(H\to G)$ and then consider $k\coloneqq|V(H)|$ as a parameter, a notion from parameterized complexity [13, 19]. In this area, fixed-parameter tractable algorithms with running time $f(k)n^{O(1)}$ for computable functions $f$ are sought. Similar to the classical notion of #P-hardness, the notion of #W[1]-hardness in parameterized complexity allows us to argue against the existence of such algorithms.
[…]
Considering fixed pattern graphs, it was shown that ETH implies a constant $\alpha>0$ such that, for every $k$-vertex graph $H$, the number #$\#\mathrm{IndSub}(H\to\,\star\,)$ cannot be computed in time $O(n^{\alpha\cdot k})$ [8, 16]. This establishes counting induced subgraphs as a hard problem, regardless of the pattern $H$.
Counting Small Induced Subgraphs with a Fixed Property
Generalizing the problem of counting induced $k$-vertex subgraphs isomorphic to a fixed graph $H$, one can fix a set $\Phi$ of $k$-vertex graphs and count induced subgraphs of $G$ that are isomorphic to some graph $H\in\Phi$. We denote this value by #$\#\mathrm{IndSub}(\Phi\to G)$. The systematic study of such counting problems was initiated by Jerrum and Meeks [26, 27, 28]: Given a fixed graph property $\Phi$, they define the parameterized problem #$\#\textnormal{{ParIndSub}}(\Phi)$ that asks, on input $(G,k)$, to determine the number #$\#\mathrm{IndSub}(\Phi^{(k)}\to G)$ of induced $k$-vertex subgraphs of $G$ that satisfy $\Phi$. Here, the slice $\Phi^{(k)}$ denotes the restriction of $\Phi$ to graphs on exactly $k$ vertices. Note that $\Phi$ itself may contain graphs of arbitrary sizes in this problem. Since their initial paper [26], it was shown that #$\#\textnormal{{ParIndSub}}(\Phi)$ is either fixed-parameter tractable or #W[1]-hard, for every computable property $\Phi$ [11]. However, the criterion on $\Phi$ for tractability of #$\#\textnormal{{ParIndSub}}(\Phi)$ remains unknown.
Clearly, if all but finitely many slices of $\Phi$ are trivial (we also say $k$-trivial) in the sense that they are empty or contain all $k$-vertex graphs, then #$\#\textnormal{{ParIndSub}}(\Phi)$ is easy to compute. Such properties $\Phi$ are called meager. For all other properties, the following was conjectured.
Conjecture 1.1 ([15, 20, 45]).
For every computable and non-meager property $\Phi$, the parameterized problem #$\#\textnormal{{ParIndSub}}(\Phi)$ is #W[1]-hard.
In the last decade, this conjecture has been verified for several special classes of non-meager properties $\Phi$. Below, we list a set of maximal results subsuming most of the known results in the area, see also [16, Section 1.1]. When we say that ETH rules out an exponent of $o(f(k))$, we mean that an algorithm with running time $n^{o(f(k))}$ for #$\#\textnormal{{ParIndSub}}(\Phi)$ would refute ETH. Even stronger, ETH implies a constant $\alpha>0$ such that #$\#\mathrm{IndSub}(\Phi^{(k)}\to G)$ cannot be computed in time $O(n^{\alpha\cdot f(k)})$ on $n$-vertex graphs, for an infinite set of integers $k$.
  1. 1.
    If $\Phi$ is closed under deleting edges (edge-monotone) [16] or closed under deleting vertices (hereditary) [20], then #$\#\textnormal{{ParIndSub}}(\Phi)$ is #W[1]-hard. This strengthens an earlier result that required $\Phi$ to satisfy both closure properties [45]. It is also shown that ETH rules out an exponent of $o(\sqrt{\log k}/\log\log k)$ for edge-monotone properties, and an exponent of $o(k)$ for hereditary properties.
  2. 2.
    Given $k\in{\mathbb{N}}$, set $\beta_{\Phi}(k)=0$ if no $k$-vertex graph satisfies $\Phi$. Otherwise, let $\beta_{\Phi}(k)$ be the number of edge-counts avoided by $\Phi$ on $k$-vertex graphs, i.e., the number of integers $\ell$ such that no graph $H$ with $k$ vertices and $\ell$ edges satisfies $\Phi$. If $\beta_{\Phi}(k)\in\omega(k)$, then #$\#\textnormal{{ParIndSub}}(\Phi)$ is #W[1]-hard [45]. Moreover, ETH rules out an exponent of $o(\beta_{\Phi}(k)/\sqrt{\log\beta_{\Phi}(k)})$.
  3. 3.
    If $\Phi$ depends only on the number of edges of the input graph, then #$\#\textnormal{{ParIndSub}}(\Phi)$ is #W[1]-hard and ETH rules out an exponent of $o(k/\sqrt{\log k})$ [45].
  4. 4.
    If there are infinitely many prime powers $t$ such that $\Phi$ holds for exactly one of $K_{t,t}$ or $I_{2t}$, then #$\#\textnormal{{ParIndSub}}(\Phi)$ is #W[1]-hard and ETH rules out an exponent of $o(k)$ [15].
In this paper, we strengthen all but the last [^1] In fact, our proofs in Section 5.2 implicitly also subsume and extend this last hardness result. of the above hardness results in several ways, e.g., by proving hardness for a larger collection of properties, by proving hardness even for modular counting variants, by proving tighter lower bounds under ETH, or by showing that the problems remain hard even when weights are incorporated into $\Phi$.
However, we consider the main contribution of our paper to be the introduction of Fourier analysis techniques to the study of #$\#\textnormal{{ParIndSub}}(\Phi)$. Indeed, the well-established toolkit of Fourier analysis allows us to derive many of the above results as relatively straightforward consequences of known results or even exercises in textbooks [39].
From Induced Subgraphs to Subgraphs
Following other recent works, we approach #$\#\textnormal{{ParIndSub}}(\Phi)$ by interpreting its $k$-vertex slices $\Phi^{(k)}$ as linear combinations of other counting functions, namely (not necessarily induced) subgraph counts: Via the inclusion-exclusion principle [33, (5.17)], it can be shown that there exist pattern graphs $H_{1},\ldots,H_{s}$, each on exactly $k$ vertices, and coefficients $\alpha_{1},\ldots,\alpha_{s}\in{\mathbb{Z}}$ such that
#$\#\mathrm{IndSub}(\Phi^{(k)}\to\,\star\,)=\sum_{i=1}^{s}\alpha_{i}\cdot\#\mathrm{Sub}(H_{i}\to\,\star\,).$ (1)
The patterns and coefficients are unique up to isomorphism, as the functions #$\#\mathrm{Sub}(H\to\,\star\,)$ are linearly independent for sets of pairwise non-isomorphic graphs.
Linear combinations as the right-hand side of (1) above enjoy a useful “complexity monotonicity” property: Under certain conditions, the presence of a single hard function #$\#\mathrm{Sub}(H\to\,\star\,)$ in the linear combination renders the evaluation of the entire linear combination hard. This phenomenon was already observed and exploited for homomorphism counts [11], for which it holds without further conditions. For subgraph counts, two conditions are required to obtain a similar statement. First, all terms in the linear combination need to feature the same number of vertices. This condition is guaranteed in the linear combinations of the form (1) we consider. Second, not only #$\#\mathrm{Sub}(H\to\,\star\,)$ is required to be hard, but even its colorful variant, where subgraph copies of $H$ containing exactly one vertex of each color in a vertex-colored graph $G$ are counted. [^2] Note that the coloring of $G$ need not be a proper coloring in the graph-theoretic sense. Such $H$-copies in $G$ are called colorful. For every fixed graph $H$, the colorful variant can be efficiently reduced to the uncolored variant; however, there exist patterns $H$ for which the colorful variant is easy while the uncolored variant is hard.
This suggests that we can understand the complexity of #$\#\mathrm{IndSub}(\Phi^{(k)}\to\,\star\,)$ by determining which patterns $H$ make counting colorful $H$-subgraph copies hard and by analyzing which patterns $H$ appear with non-zero coefficients in the linear combination (1) derived from $\Phi$. Similar approaches were taken in [15, 16, 45].
Which subgraph patterns are hard to count?
The complexity of counting subgraphs is well-understood in both the uncolored and colorful settings [12, 11]: First, observe that easy patterns emerge when counting subgraphs rather than induced subgraphs. For example, the number of uncolored subgraph copies of the edgeless $k$-vertex graph $I_{k}$ in an $n$-vertex graph $G$ is simply $\binom{n}{k}$. If $G$ has $k$ color classes of sizes $n_{1},\ldots,n_{k}$, then the number of colorful subgraph copies of $I_{k}$ in $G$ equals $\prod_{i}n_{i}$. Note however that counting induced copies of $I_{k}$ in $G$ amounts to the hard problem of counting independent sets in $G$. Similarly, counting colorful independent sets is also hard.
In the uncolored setting, it is known that the subgraph variant #$\#\textnormal{{ParSub}}(\mathcal{H})$ of the induced subgraph counting problem #$\#\textnormal{{ParIndSub}}(\mathcal{H})$ is polynomial-time solvable if the maximum vertex-cover number among graphs in $\mathcal{H}$ is finite. Otherwise, the problem is #W[1]-hard under the technical provision that $\mathcal{H}$ is recursively enumerable, and almost tight lower bounds are known under ETH [12, 11].
The colorful variant, which is relevant in this paper, admits additional tractable patterns: A standard dynamic programming approach yields an $O(n^{\operatorname{tw}(H)+1})$ time algorithm for counting colorful subgraphs isomorphic to $H$ in an $n$-vertex graph $G$ with vertex-colors. Here, $\operatorname{tw}(H)$ is the treewidth of $H$, a graph parameter that measures how similar $H$ is to a forest for algorithmic purposes. In particular, there are graphs $H$ of large vertex-cover number and treewidth $1$, such as matchings or paths. This implies that the problem #$\#\textnormal{{ParColSub}}(\mathcal{H})$ of counting colorful $H$-subgraph copies in a vertex-colored graph $G$, on input $H\in\mathcal{H}$ and $G$, is polynomial-time solvable if there is a finite bound on the treewidth of graphs in $\mathcal{H}$. Otherwise, the problem is #W[1]-hard, again assuming that $\mathcal{H}$ is recursively enumerable, see [12, Theorem 3.4] or [10, Theorem 5.6]. Moreover, ETH implies a constant $\beta>0$ such that for every graph $H$ with treewidth $t$, the number of colorful $H$-subgraph copies in $n$-vertex graphs cannot be computed in time $O(n^{\beta\cdot t/\log t})$ [11, 16].
In this paper, we rely on large edge-density of $H$ as a sufficient proxy condition to guarantee large treewidth of $H$ and thus hardness of counting colorful $H$-subgraph copies. Indeed, it is well-known that a graph $H$ of treewidth $t$ has at most $t\cdot|V(H)|$ edges. In a nutshell, we obtain that counting patterns $H$ with many edges is hard:
Fact 1.2.
Assuming ETH, there is a constant $\beta>0$ such that, for every pattern $H$ with $|E(H)|\geq t\cdot|V(H)|$, counting colorful $H$-subgraph copies in vertex-colored graphs with $n$ vertices requires $O(n^{\beta\cdot t/\log t})$ time.
A similar version of this fact holds with regard to #W[1]-hardness. We stress that large edge-density is only a sufficient criterion for hardness. Indeed, constant-degree expanders provide an example of patterns $H$ that have degree $3$ and treewidth $\Omega(|V(H)|)$. Nevertheless, Fact 1.2 and variants thereof are sufficient for all hardness results obtained in this paper.
Which subgraph patterns appear in the linear combination?
In view of Fact 1.2, we can establish hardness of #$\#\mathrm{IndSub}(\Phi^{(k)}\to\,\star\,)$ by finding subgraph patterns $H$ with a superlinear amount of edges in the linear combinations (1) derived from $\Phi$. This task however turns out to be nontrivial. On one hand, one can show straightforwardly that the coefficient of $H$ is given by the alternating enumerator of $\Phi$ on $H$. We define this quantity, with an extra factor of $(-1)^{|E(H)|}$ compared to previous works [43, 44, 15], as
$\widehat{\Phi}(H)=(-1)^{|E(H)|}\sum_{S\subseteq E(H)}(-1)^{|S|}\cdot\Phi(H[S]),$
where $H[S]$ is the subgraph obtained from $H$ by keeping all vertices and only the edges in $S$.
On the other hand, despite the inconspicuously explicit form of $\widehat{\Phi}(H)$, it is highly nontrivial to determine for which properties $\Phi$ and graphs $H$ we have $\widehat{\Phi}(H)\neq 0$. With some exceptions [20, 45], most approaches towards this can be traced back to a seminal paper [44] that interprets edge-monotone properties $\Phi$ defined on $k$-vertex graphs as abstract simplicial complexes. Up to sign, $\widehat{\Phi}(K_{k})$ then is the (reduced) Euler characteristic of the simplicial complex: Viewing cardinality-$t$ sets as dimension-$t$ facets of the complex corresponding to $\Phi$, we see that $\widehat{\Phi}(K_{k})$ is the difference between the counts of odd-dimension and even-dimension facets, similar to Euler’s formula for plane graphs. By understanding group actions on simplicial complexes that preserve the Euler characteristic modulo primes, it was then shown that $\widehat{\Phi}(K_{k})\neq 0$ for several properties $\Phi$, e.g., for edge-monotone properties $\Phi$ that are false on all odd cycles. In subsequent works [15, 16, 46], transitive group actions on graphs were used to derive more general results, also for related pattern counting problems, culminating in the recent result on edge-monotone properties [16] mentioned above.
Our Approach: Connecting the Subgraph Basis and Fourier Analysis
In a nutshell, we observe that Fourier analysis enables a more straightforward and versatile approach to alternating enumerators than previous methods. As mentioned above, a linear combination $f(\star)=\sum_{H}\alpha_{H}\,\#\mathrm{IndSub}(H\to\,\star\,)$ with $H$ ranging over $k$-vertex graphs also admits a subgraph count representation $f(\star)=\sum_{H}\beta_{H}\,\#\mathrm{Sub}(H\to\,\star\,)$, where $H$ also ranges over $k$-vertex graphs. Both representations are unique, as the underlying basis functions are linearly independent. We observe that the basis change between these representations is “essentially” a Fourier transform.
The precise correspondence can be phrased via polynomials. Let $X_{t}$ be the undirected complete $t$-vertex graph with an indeterminate $x_{ij}$ on edge $ij$. We define the weighted subgraph count from $H$ into $X$ as the polynomial
#$\#\mathrm{Sub}(H\to X_{t})=\sum_{F}\prod_{ij\in E(F)}x_{ij},$ (2)
where $F$ ranges over all subgraphs of $X_{t}$ isomorphic to $H$, ignoring edge-weights. Also define
#$\#\mathrm{IndSub}(H\to X_{t})=\sum_{F}\prod_{ij\in E(F)}x_{ij}\prod_{ij\in\overline{E(F)}}(1-x_{ij}),$ (3)
with $\overline{E(F)}=\binom{V(F)}{2}\setminus E(F)$. For every $t$-vertex graph $G$, we have #$\#\mathrm{Sub}(H\to G)=\#\mathrm{Sub}(H\to X_{t})$ after substituting $1$ into $x_{ij}$ for $ij\in E(G)$ and 00 for $ij\notin E(G)$, and the same holds for #$\#\mathrm{IndSub}(H\to X_{t})$. In other words, the evaluations of the polynomials on 00-$1$-inputs corresponding to $t$-vertex graphs $G$ are precisely the counts of (induced) $H$-copies in $G$. As multilinear polynomials are determined by their evaluations on 00-$1$-inputs, it follows that #$\#\mathrm{Sub}(H\to X_{t})$ and #$\#\mathrm{IndSub}(H\to X_{t})$ are the unique multilinear polynomials that agree with #$\#\mathrm{Sub}(H\to\,\star\,)$ and #$\#\mathrm{IndSub}(H\to\,\star\,)$ on $t$-vertex graphs, when the latter two functions are viewed as functions $\{0,1\}^{m}\to{\mathbb{R}}$ for $m=\binom{t}{2}$.
The key observation is now that the degrees of these polynomials give very useful insights into the complexity of #$\#\mathrm{Sub}(H\to\,\star\,)$ and linear combinations thereof. Recall from Fact 1.2 that #$\#\mathrm{Sub}(H\to\,\star\,)$ and its colorful variant are hard if $H$ has a large number of edges. The number of edges in a $k$-vertex graph $H$ however is precisely the degree of #$\#\mathrm{Sub}(H\to X_{t})$ for $t\geq k$; it suffices to choose $t=k$ in the following. Also recall the complexity monotonicity property for subgraph counts: Linear combinations $f(\star)=\sum_{H}\beta_{H}\,\#\mathrm{Sub}(H\to\,\star\,)$, with $H$ ranging over $k$-vertex graphs, are hard if they contain a function #$\#\mathrm{Sub}(H\to\,\star\,)$ for large $|E(H)|$. This in turn occurs if and only if the polynomial representation $f(X_{k})$ has large degree. In summary, one way to prove that linear combinations $f(\star)$ of subgraph counts from $k$-vertex patterns are difficult is by showing that the polynomial $f(X_{k})$ has large degree; indeed, showing a superlinear degree $\omega(k)$ is sufficient.
As an example, this reasoning shows that #$\#\mathrm{IndSub}(H\to\,\star\,)$ is hard for every $k$-vertex graph $H$, since its polynomial representation (3) clearly has degree $\binom{k}{2}$ and its representation over the subgraph count basis therefore contains a graph on $\binom{k}{2}$ edges, i.e., the complete graph $K_{k}$. Taking linear combinations of the functions #$\#\mathrm{IndSub}(H\to\,\star\,)$ for different $H$ may however lead to cancellations that make $K_{k}$ disappear from this representation.
To understand the degree in the representing polynomial for more general functions $f$, the rich toolkit of Fourier analysis is available [29, 39]. In a nutshell, the degrees we wish to understand are precisely the Fourier degrees of Boolean functions, and various well-developed techniques are available to obtain lower bounds on them. Recall that, in the polynomial representations above, the function of interest is interpreted as $f:\{0,1\}^{m}\to{\mathbb{R}}$ for $m=\binom{k}{2}$, and the degree of the unique multilinear polynomial agreeing with $f$ determined the complexity of $f$. In Fourier analysis, we view $f:\{-1,1\}^{m}\to{\mathbb{R}}$ by interpreting 00 as $1$ and $1$ as $-1$. The unique multilinear polynomial agreeing with $f$ is called the Fourier expansion of $f$, and its degree is the Fourier degree of $f$. While the different representations of inputs lead to different representing multilinear polynomials, they can be transformed into each other by a simple affine change of variables, implying that they have the same degree. We can thus establish that linear combinations $f(\star)$ of subgraph counts from $k$-vertex patterns are hard by showing that they have large Fourier degree, interpreted as $f:\{-1,1\}^{m}\to{\mathbb{R}}$.
Our Results
Sparse and Hereditary Properties.
Our first result applies the Fourier machinery to obtain hardness of moderately sparse properties, i.e., properties that are missing an exponential fraction of graphs. This implies known results on hereditary properties [20] as a special case. Moreover, our result holds even for “properties” that do not output 00 or $1$. In the following, a graph invariant $\Phi$ is an isomorphism-invariant function from graphs to ${\mathbb{Q}}$. If there is some $k\in{\mathbb{N}}$ such that all graphs $G$ with $\Phi(G)\neq 0$ have exactly $k$ vertices, then we say that $\Phi$ is a $k$-vertex graph invariant and define
#$\#\mathrm{IndSub}(\Phi\to G)=\sum_{F}\Phi(F)\cdot\#\mathrm{IndSub}(F\to G).$
The uncertainty principle establishes that every function $f\colon\{-1,1\}^{m}\to{\mathbb{R}}$ with small support contains many terms in its Fourier expansion and thus has large Fourier degree [39]. More formally, for every $10<\varepsilon<$ 1 there is some $\delta>0$ such that, if $1\leq|\operatorname{supp}(f)|\leq(2-\varepsilon)^{m}$, then the Fourier degree of $f$ is at least $\delta m$. Using this, we show that the subgraph basis representation of #$\#\mathrm{IndSub}(\Phi\to\star)$ contains a graph $H$ with $\delta\cdot\binom{k}{2}$ edges, rendering #$\#\mathrm{IndSub}(\Phi\to\star)$ hard.
Theorem 1.3.
For every $10<\varepsilon<$ 1 there is $\delta>0$ such that, for every $k$-vertex graph invariant $\Phi$ with $1\leq|\operatorname{supp}(\Phi)|\leq(2-\varepsilon)^{\binom{k}{2}}$, no algorithm computes #$\#\mathrm{IndSub}(\Phi\to G)$ in time $O(n^{\delta\cdot k/\sqrt{\log k}})$ on $n$-vertex graphs $G$ unless ETH fails.
Moreover, for every computable graph invariant $\Phi$ such that $1\leq|\operatorname{supp}(\Phi^{(k)})|\leq(2-\varepsilon)^{\binom{k}{2}}$ holds for infinitely many slices $\Phi^{(k)}$, the problem #$\#\textnormal{{ParIndSub}}(\Phi)$ is #W[1]-hard. This holds in particular when $\Phi$ is a hereditary property that is not meager.
The consequence for hereditary properties follows from the fact that every non-empty property excluding a fixed induced subgraph $F$ contains at most $(2-\varepsilon)^{\binom{k}{2}}$ graphs in the $k$-th slice if $k$ is sufficiently large, depending on $F$ [40].
(Edge-)Monotone Properties & Modular Counting.
Let $\Phi$ be a nontrivial $k$-vertex graph property that is monotone, i.e., closed under the deletion of edges. [^3] In the literature, this condition is often called “edge-monotone” (see e.g., [16, 20, 45]). By adapting known results on the Fourier degrees of monotone functions with symmetry conditions [14, 41], we show that the subgraph expansion of $\Phi$ contains some $k$-vertex graph $H$ containing as subgraph the complete bipartite graph $K_{\ell,\ell}$ for some $\ell>k/4$. This gives us particularly strong bounds under ETH and holds even when the subgraph expansion is reduced modulo a prime $p$; then we obtain $\ell>k/p^{2}$.
Theorem 1.4.
There is $\delta>0$ such that, for every $k$-vertex graph property that is monotone and not $k$-trivial, no algorithm computes #$\#\mathrm{IndSub}(\Phi\to G)$ in time $O(n^{\delta\cdot k})$ on $n$-vertex graphs $G$ unless ETH fails. Moreover, for every computable graph property $\Phi$ that, for infinitely many $k$, is monotone on $k$-vertex graphs and not $k$-trivial, the problem #$\#\textnormal{{ParIndSub}}(\Phi)$ is #W[1]-hard.
For every fixed prime $p$, both statements hold even when the value #$\#\mathrm{IndSub}(\Phi\to G)$ is only to be computed modulo $p$, assuming the randomized ETH instead of ETH.
Fully Symmetric Properties.
For fully symmetric properties, i.e., properties that depend only on the number of edges, we obtain strong hardness by straightforward applications of known Fourier analysis results for fully symmetric functions. Special cases of such properties were studied before [26, 28, 44], and [45] shows that non-trivial and fully symmetric graph properties $\Phi$ induce hard problems #$\#\textnormal{{ParIndSub}}(\Phi)$. However, [45] only rules out algorithms with running time $n^{o(k/\sqrt{\log k})}$ assuming ETH. We obtain tight lower bounds under ETH with an arguably simpler proof.
Theorem 1.5.
There is $\delta>0$ such that, for every $k$-vertex graph property $\Phi$ that is fully symmetric and not $k$-trivial, no algorithm computes #$\#\mathrm{IndSub}(\Phi\to G)$ in time $O(n^{\delta\cdot k})$ on $n$-vertex graphs $G$ unless ETH fails.
Moreover, for every computable graph property $\Phi$ that is fully symmetric on $k$-vertex graphs and not $k$-trivial for infinitely many $k$, the problem #$\#\textnormal{{ParIndSub}}(\Phi)$ is #W[1]-hard.
Properties Avoiding many Hamming Weights.
We also obtain hardness for properties $\Phi$ that avoid a superlinear number of Hamming weights, a setting that was also studied before [26, 45]. Given a $k$-vertex graph invariant $\Phi$, we define its Hamming weight set as ${\sf hw}(\Phi)\coloneqq\{|E(H)|\mid H\in\Phi\}$, i.e., the set of all edge-counts among the graphs in $\Phi$. We then define a value $\beta_{\Phi}$ that equals 00 when $\Phi$ is empty and $\binom{k}{2}-|{\sf hw}(\Phi)|$ otherwise. For a general graph invariant $\Phi$ and $k\in{\mathbb{N}}$, we define $\beta_{\Phi}(k)\coloneqq\beta_{\Phi^{(k)}}$. By a symmetrization argument similar to [38], we obtain:
Theorem 1.6.
There is $\delta>0$ such that, for every $20<d\leq k/$ 2 and $k$-vertex graph invariant $\Phi$ with $\beta_{\Phi}\geq dk,$ no algorithm computes #$\#\mathrm{IndSub}(\Phi\to G)$ in time $O(n^{\delta\cdot d/\sqrt{\log d}})$ on $n$-vertex graphs $G$ unless ETH fails. Moreover, #$\#\textnormal{{ParIndSub}}(\Phi)$ is #W[1]-hard for computable graph invariants $\Phi$ with $\beta_{\Phi}(k)\in\omega(k)$.
The All-Even Property.
Finally, consider the curious property $\Phi_{\sf even}$ that contains exactly those graphs where all vertices have even degree. This property turns out to be interesting, since only few graphs $H$ are contained in its subgraph expansion; we can identify them as precisely the bipartite graphs. This implies that #$\#\textnormal{{ParIndSub}}(\Phi_{\sf even})$ is #W[1]-hard, with tight lower bounds under ETH, but it also yields non-trivial upper bounds on its complexity. Moreover, the fact that only bipartite graphs appear in the subgraph expansion is actually useful to prove hardness of other properties, as we can take pointwise products with $\Phi_{\sf even}$ to filter out unwanted graphs from a property $\Phi$ and translate degree lower bounds and complexity implications for the filtered version back to $\Phi$. In particular, we can obtain the following variant of Theorem 1.3.
Theorem 1.7.
For every $10<\varepsilon<$ 1 there is $\delta>0$ such that, for every $k$-vertex graph invariant $\Phi$ with $1\leq|\operatorname{supp}(\Phi\cdot\Phi_{\sf even})|\leq(\sqrt{2}-\varepsilon)^{\binom{k}{2}}$, no algorithm computes #$\#\mathrm{IndSub}(\Phi\to G)$ in time $O(n^{\delta\cdot k/\sqrt{\log k}})$ on $n$-vertex graphs $G$ unless ETH fails.
Moreover, for every computable graph invariant $\Phi$ such that $1\leq|\operatorname{supp}(\Phi^{(k)}\cdot\Phi_{\sf even})|\leq(\sqrt{2}-\varepsilon)^{\binom{k}{2}}$ holds for infinitely many slices $\Phi^{(k)}$, the problem #$\#\textnormal{{ParIndSub}}(\Phi)$ is #W[1]-hard.
Here, $\Phi\cdot\Phi_{\sf even}$ denotes the point-wise product of $\Phi$ and $\Phi_{\sf even}$ defined via $(\Phi\cdot\Phi_{\sf even})(G)=\Phi(G)\cdot\Phi_{\sf even}(G)$. In particular, $\operatorname{supp}(\Phi\cdot\Phi_{\sf even})$ is the intersection of the supports of $\Phi$ and $\Phi_{\sf even}$.
Weisfeiler-Leman Dimension of Induced Subgraph Counts
Our techniques also allow us to bound or even precisely determine the Weisfeiler-Leman dimension of graph parameters. The $k$-dimensional Weisfeiler-Leman algorithm ($k$-WL) is a standard heuristic for testing isomorphism of graphs with connections to various other areas [23, 24, 30, 36]. The Weisfeiler-Leman dimension of a graph invariant $f$ is the minimal $k\geq 1$ such that $k$-WL distinguishes between all graphs $G,G^{\prime}$ for which $f(G)\neq f(G^{\prime})$. A high value of this complexity measure implies inexpressibility in certain logics and lower bounds for symmetric computational models.
Our results yield bounds on the WL-dimension of #$\#\mathrm{IndSub}(\Phi\to\,\star\,)$ for various $k$-vertex graph invariants $\Phi$. For example, for monotone $k$-vertex graph properties $\Phi$, the WL-dimension of #$\#\mathrm{IndSub}(\Phi\to\,\star\,)$ is $>\frac{k}{4}$. The WL-dimension of #$\#\mathrm{IndSub}(\Phi_{\sf even}^{(k)}\to\,\star\,)$ is equal to $\left\lfloor k/2\right\rfloor$.
Organization of the Paper
In Section 2, we give formal definitions of the graph-theoretic concepts used, brief introductions to parameterized and exponential-time complexity, and a synopsis of basic concepts from Fourier analysis. We describe the basis transformation between induced subgraph counts and subgraph counts and complexity results for the latter basis in Section 3 before connecting this transformation to Fourier analysis in Section 4. Based on this foundation, we then prove our main hardness results in Section 5 and the results on the all-even property in Section 6. Finally, we adapt the proofs to obtain lower bounds for the Weisfeiler-Leman dimension of induced subgraph counts in Section 7. The conclusions list several open problems, and the appendix contains technical proofs omitted from the main paper to improve reading flow.
2 Preliminaries
We write ${\mathbb{Z}}$ for the integers and ${\mathbb{Z}}_{p}=\{0,1,\dots,p-1\}$ for the ring with addition and multiplication taken modulo $p$. For an integer $n\geq 1$, we define $[n]\coloneqq\{1,\dots,n\}$. For a set $X$ and an integer $k\geq 0$, we use $\binom{X}{k}$ to denote the collection of $k$-element subsets of $X$.
2.1 Graphs
We use standard graph notation. All graphs considered in this paper are undirected and simple, i.e., they contain no multiedges or self-loops. We write $V(G)$ and $E(G)$ to denote the vertex set and edge set of a graph $G$, respectively. Also, we use $vw$ as a shorthand for an edge $\{v,w\}\in E(G)$.
A graph $H$ is a subgraph of another graph $G$, denoted $H\subseteq G$, if $V(H)\subseteq V(G)$ and $E(H)\subseteq E(G)$. For a graph $G$ and a set $X\subseteq V(G)$, we write $G[X]=(X,\{vw\in E(G)\mid v,w\in X\})$ to denote the (vertex-)induced subgraph of $G$ by $X$. Similarly, for $S\subseteq E(G)$, we write $G[S]=(V(G),S)$ to denote the (edge-)induced subgraph of $G$ by $S$. It is usually clear from the context whether we are taking a vertex- or an edge-induced subgraph.
For $k\geq 1$, we write $K_{k}$ for the complete graph with $k$ vertices, i.e., $V(K_{k})=[k]$ and $E(K_{k})=\binom{[k]}{2}$. For $k,\ell\geq 1$, we write $K_{k,\ell}$ to denote the complete bipartite graph with $k+\ell$ vertices, i.e., $V(K_{k,\ell})=[k+\ell]$ and $E(K_{k,\ell})=\{\{i,k+j\}\mid i\in[k],j\in[\ell]\}$. We remark that we explicitly define the graph $K_{k}$ to have vertex set $[k]$. In particular, for $S\subseteq\binom{[k]}{2}$, we commonly use $K_{k}[S]$ to refer to the graph with vertex set $[k]$ and edge set $S$. If the index $k$ is clear from context, we sometimes omit it and simply write $K[S]$.
A graph invariant $\Phi$ (over a field ${\mathbb{F}}$) maps each graph $G$ to a value $\Phi(G)\in{\mathbb{F}}$ such that $\Phi(G)=\Phi(G^{\prime})$ for all isomorphic graphs $G,G^{\prime}$. In this work, we only consider graph invariants over the rationals ${\mathbb{Q}}$ and over ${\mathbb{Z}}_{p}$ for prime $p$. Unless stated otherwise, graph invariants are defined over the rationals. A graph invariant is a graph property if $\Phi(G)\in\{0,1\}$ for all graphs $G$. For a graph $G$ and a graph property $\Phi$, we write $G\in\Phi$ as a shorthand for $\Phi(G)=1$. For $k\geq 1$, a $k$-vertex graph invariant/property is a graph invariant/property $\Phi$ such that $\Phi(G)=0$ if $|V(G)|\neq k$. For a graph invariant $\Phi$, we write $\Phi^{(k)}$ to denote the graph invariant on $k$-vertex graphs derived by $\Phi$, i.e.,
$\Phi^{(k)}(G)=\begin{cases}\Phi(G)&\text{if $|V(G)|=k$}\\ 0&\text{otherwise}\end{cases}.$
For $k\geq 1$, a graph invariant $\Phi$ (over ${\mathbb{F}}$) is $k$-trivial if there is some $c\in{\mathbb{F}}$ such that $\Phi(G)=c$ for all $k$-vertex graphs $G$. A graph property $\Phi$ is meager if it is $k$-trivial for all but finitely many $k$.
For a graph invariant $\Phi$, we define
#$\#\mathrm{IndSub}(\Phi\to G)\coloneqq\sum_{X\subseteq V(G)}\Phi(G[X]).$
Observe that, if $\Phi$ is a graph property, then #$\#\mathrm{IndSub}(\Phi\to G)$ is the number of vertex-induced subgraphs of $G$ that satisfy $\Phi$. We write #$\#\mathrm{IndSub}(\Phi\to\,\star\,)$ for the map $G\mapsto\#\mathrm{IndSub}(\Phi\to G)$; this is again a graph invariant (over the same field as $\Phi$).
2.2 Parameterized Complexity
We give a brief introduction to parameterized counting complexity and refer the reader to the textbooks [13, 19] and theses [10, 43] for additional background.
A parameterized counting problem consists of a function $P\colon\Sigma^{*}\to{\mathbb{Q}}$ and a parameterization $\kappa\colon\Sigma^{*}\to{\mathbb{Z}}_{\geq 0}$. For an instance $x\in\Sigma^{*}$, the number $\kappa(x)$ is called the parameter of $x$. The problem $(P,\kappa)$ is called fixed-parameter tractable (fpt) if there is a computable function $f$, a constant $c\in{\mathbb{Z}}_{\geq 1}$ and a (deterministic) algorithm ${\mathcal{A}}$ that, given an instance $x\in\Sigma^{*}$, computes $P(x)$ in time $f(\kappa(x))\cdot|x|^{c}$.
[…]
For a graph invariant $\Phi$, the parameterized counting problem #$\#\textnormal{{ParIndSub}}(\Phi)$ takes as input a graph $G$ and an integer $k\geq 1$ (the parameter) and outputs #$\#\mathrm{IndSub}(\Phi^{(k)}\to G)$. When studying the complexity of #$\#\textnormal{{ParIndSub}}(\Phi)$, it will mostly suffice to restrict our attention to the $k$-th slice of the problem, for fixed integers $k\geq 1$, where the input consists only of a graph $G$. This motivates us to define the following problem: Given $k\geq 1$ and a $k$-vertex graph invariant $\Phi$, the problem #$\#\textnormal{{IndSub}}(\Phi)$ takes as input a graph $G$ and outputs #$\#\mathrm{IndSub}(\Phi\to G)$. For a graph invariant $\Phi$, we also write #$\#\textnormal{{IndSub}}(\Phi,k)$ instead of #$\#\textnormal{{IndSub}}(\Phi^{(k)})$. Compared to #$\#\textnormal{{ParIndSub}}(\Phi)$, this problem is polynomial-time solvable for any fixed $\Phi$ and $k$.
Modular Counting.
[…]
For counting induced subgraphs, we only consider modular counting versions for graph properties $\Phi$. In particular, if $\Phi$ is a graph property, then #$\#\mathrm{IndSub}(\Phi\to G)$ is an integer and we can safely define #$\#_{p}\mathrm{IndSub}(\Phi\to G)\in\{0,\dots,p-1\}$. In turn, this allows us to define the problems #$\#_{p}\textnormal{{ParIndSub}}(\Phi)$ and #$\#_{p}\textnormal{{IndSub}}(\Phi)$ for graph properties $\Phi$.
Fine-Grained Lower Bounds.
To obtain more fine-grained lower bounds on the complexity of #$\#\textnormal{{ParIndSub}}(\Phi)$ as well as #$\#\textnormal{{IndSub}}(\Phi)$, we also rely on the Exponential Time Hypothesis (ETH) dating back to [25] (see, e.g., [13, Conjecture 14.1]).
Conjecture 2.1 (Exponential Time Hypothesis (ETH)).
There is a real number $\varepsilon>0$ such that the problem $3$-Sat cannot be solved in time $O(2^{\varepsilon n})$, where $n$ denotes the number of variables of the input formula.
In the setting of modular counting, we rely on the following slightly stronger conjecture that even rules out randomized algorithms.
Conjecture 2.2 (Randomized Exponential Time Hypothesis (rETH)).
There is a real number $\varepsilon>0$ such that the problem $3$-Sat cannot be solved in time $O(2^{\varepsilon n})$ by a randomized algorithm with error probability at most $1/3$, where $n$ denotes the number of variables of the input formula.
2.3 Boolean Functions
Let $f\colon\{-1,1\}^{m}\to{\mathbb{R}}$ be a Boolean function of arity $m$. The reader is encouraged to associate $-1$ with True and $1$ with False. The Fourier expansion of $f$ is its unique representation as a multilinear polynomial in ${\mathbb{R}}[x_{1},\ldots,x_{m}]$. For $S\subseteq[m]$, define the monomial corresponding to $S$ as
$x^{S}\coloneqq\prod_{i\in S}x_{i}.$
Then there are unique so-called Fourier coefficients $\widehat{f}(S)\in{\mathbb{R}}$ for $S\subseteq[m]$ such that
$f(\boldsymbol{x})=\sum_{S\subseteq[m]}\widehat{f}(S)x^{S}.$ (4)
See [39, Theorem 1.1] for a proof of the existence and uniqueness of (4), and the entire textbook [39] for extensive background on the Fourier expansion of Boolean functions. We write
$\operatorname{supp}(f)$$\coloneqq\{\boldsymbol{x}\in\{-1,1\}^{m}\mid f(\boldsymbol{x})\neq 0\},$
$\operatorname{supp}(\widehat{f})$$\coloneqq\{S\subseteq[m]\mid\widehat{f}(S)\neq 0\}.$
By the uncertainty principle, one of these two sets is always large; see [39, Exercise 3.15].
Lemma 2.3 (Uncertainty Principle).
If $f\colon\{-1,1\}^{m}\to{\mathbb{R}}$ is not the zero function, then
$|\operatorname{supp}(f)|\cdot|\operatorname{supp}(\widehat{f})|\geq 2^{m}.$
The degree $\deg(f)$ of $f$ is the degree of the Fourier expansion, i.e.,
$\deg(f)\coloneqq\max\{|S|\mid S\in\operatorname{supp}(\widehat{f})\}.$
By standard upper bounds on binomial coefficients (see, e.g., [19, Lemma 16.19]), small degree of $f$ implies small support:
Lemma 2.4.
Let $\mathbf{H}(c)=-c\log_{2}(c)-(1-c)\log_{2}(1-c)$ be the binary entropy function. For every function $f\colon\{-1,1\}^{m}\to{\mathbb{R}}$ with $\deg(f)\leq cm$ for $20<c\leq 1/$ 2, we have
[…]
Conversely, if $|\operatorname{supp}(\widehat{f})|>2^{\mathbf{H}(c)m}$, then $\deg(f)>cm$.
We regularly need to restrict Boolean functions to subcubes. Let $f\colon\{-1,1\}^{m}\to{\mathbb{R}}$ be a Boolean function of arity $m$ and let $J\subseteq[m]$. We write $\overline{J}\coloneqq[m]\setminus J$ for the complement of $J$ in $[m]$. Also let $\boldsymbol{z}=(z_{j})_{j\in\overline{J}}\in\{-1,1\}^{\overline{J}}$. Then we define the restriction of $f$ to $J$ using $\boldsymbol{z}$, denoted by $f_{J|\boldsymbol{z}}\colon\{-1,1\}^{J}\to{\mathbb{R}}$, to be subfunction of $f$ given by fixing the coordinates in $\overline{J}$ to the values in $\boldsymbol{z}$. Slightly abusing notation, for $\boldsymbol{y}\in\{-1,1\}^{J}$, we write $(\boldsymbol{y},\boldsymbol{z})$ for the composite tuple in $\{-1,1\}^{m}$, even tough $\boldsymbol{y}$ and $\boldsymbol{z}$ are not actually concatenated. In this notation, we have $f_{J|\boldsymbol{z}}(\boldsymbol{y})=f(\boldsymbol{y},\boldsymbol{z})$. Consistent with previous notation, we use $\widehat{f}_{J|\boldsymbol{z}}(S)\in{\mathbb{R}}$ for $S\subseteq J$ to denote the Fourier coefficients of $f_{J|\boldsymbol{z}}$. Then
$\widehat{f}_{J|\boldsymbol{z}}(S)=\sum_{T\subseteq\overline{J}}\widehat{f}(S\cup T)\prod_{j\in T}z_{j}$ (5)
for all $S\subseteq J$; see [39, Proposition 3.21]. In particular, $\deg(f)\geq\deg(f_{J|\boldsymbol{z}})$.
3 Complexity in the Subgraph Basis
It is known that pattern counting problems can be expressed and studied in different bases. For example, the problems studied in this paper admit direct expressions as linear combinations of induced subgraph counts for graphs $F_{1},\ldots,F_{s}$ and coefficients $\alpha_{1},\ldots,\alpha_{s}\in{\mathbb{Q}}$, i.e.,
$f(\,\star\,)=\sum_{i=1}^{s}\alpha_{i}\cdot\#\mathrm{IndSub}(F_{i}\to\,\star\,).$ (6)
Such problems can also be expressed as linear combinations of not necessarily induced subgraph counts for graphs $H_{1},\ldots,H_{s}$ and coefficients $\beta_{1},\ldots,\beta_{s}\in{\mathbb{Q}}$, i.e.,
$f(\,\star\,)=\sum_{j=1}^{t}\beta_{j}\cdot\#\mathrm{Sub}(H_{j}\to\,\star\,).$ (7)
See [33, (5.17)] for a proof. The subgraph basis representation (7) turns out to be more useful for understanding the complexity of $f$. In the transformation from (6) to (7), and in this paper in general, the following notion is central:
Definition 3.1.
The alternating enumerator of a graph invariant $\Phi$ on a graph $H$ is
$\widehat{\Phi}(H)=(-1)^{|E(H)|}\sum_{S\subseteq E(H)}(-1)^{|S|}\,\Phi(H[S]).$
Remark 3.2.
In earlier works [15, 16, 44], the alternating enumerator is defined without the factor $(-1)^{|E(H)|}$. These works aim at showing that $\widehat{\Phi}(H)\neq 0$ for certain graphs $H$; such statements are not affected by the absence of the factor $(-1)^{|E(H)|}$. To obtain more straightforward statements in Lemmas 3.3, 4.1 and 5.19, we prefer the variant defined above.
The coefficients in the subgraph representation of a graph invariant can be shown to be alternating enumerators. While similar correspondences are known, see [15, Lemma 8] or [44, Theorem 1], we could not find the precise lemma below in the literature. The appendix includes an explicit proof based on the inclusion–exclusion principle.
Lemma 3.3.
Given a $k$-vertex graph invariant $\Phi$, for $k\geq 1$, we have
#$\#\mathrm{IndSub}(\Phi\to\star)=\sum_{H}\widehat{\Phi}(H)\cdot\#\mathrm{Sub}(H\to\star),$ (8)
where $H$ ranges over all unlabelled $k$-vertex graphs.
To show hardness of the problem #$\#\textnormal{{IndSub}}(\Phi)$ (and its variants) we follow previous work [16, 15, 44] and identify graphs of unbounded treewidth with non-zero alternating enumerator: For a graph $H$, we write $\operatorname{tw}(H)$ for the treewidth of $H$ (see, e.g., [13, Chapter 7]) and say that an (infinite) sequence of graphs $(H_{k})_{k\geq 1}$ has unbounded treewidth if, for every $t\geq 1$, there is some $k\geq 1$ such that $\operatorname{tw}(H_{k})\geq t$.
Theorem 3.4.
Let $\Phi$ be a computable graph invariant and let $(H_{k})_{k\geq 1}$ be a sequence of graphs of unbounded treewidth.
  1. (a)
    If $\widehat{\Phi}(H_{k})\neq 0$ for all $k\geq 0$, then #$\#\textnormal{{ParIndSub}}(\Phi)$ is #W[1]-hard.
  2. (b)
[…]
Also, we use the next theorem to obtain more fine-grained lower bounds based on ETH and rETH.
Theorem 3.5.
There is a universal constant $\alpha_{\textsc{ind}}>0$ and an integer $N_{0}\geq 1$ such that for all numbers $k,\ell\geq 1$, the following holds:
  1. (a)
    If $\Phi$ is a $k$-vertex graph invariant and there exists a graph $H$ with $\widehat{\Phi}(H)\neq 0$ and $E(H)\geq k\cdot\ell\geq N_{0}$, then #$\#\textnormal{{IndSub}}(\Phi)$ cannot be solved in time $O(n^{\alpha_{\textsc{ind}}\cdot\ell/\sqrt{\log\ell}})$ unless ETH fails.
  2. (b)
    Under the same conditions as above, if $H$ is additionally a supergraph of $K_{\ell,\ell}$, then #$\#\textnormal{{IndSub}}(\Phi)$ cannot be solved in time $O(n^{\alpha_{\textsc{ind}}\cdot\ell})$ unless ETH fails.
  3. (c)
    If $p$ is a prime and the condition $\widehat{\Phi}(H)\neq 0$ is replaced by $\widehat{\Phi}(H)\neq 0\bmod p$, then Items (a) and (b) hold for #$\#_{p}\textnormal{{ParIndSub}}(\Phi)$ instead of #$\#\textnormal{{ParIndSub}}(\Phi)$, assuming rETH instead of ETH.
Both theorems can be proved by (variants of) known arguments [10, 15, 16, 44] and several special cases have already been stated in the literature (see, e.g., [16]). Still, we provide full proofs in Appendix A.
Finally, let us also note the following algorithmic result. Let $G$ be a graph. A vertex cover of $G$ is a set $C\subseteq V(G)$ such that $e\cap C\neq\emptyset$ for every edge $e\in E(G)$. The vertex cover number of $G$, denoted by $\operatorname{vc}(G)$, is the minimal size of a vertex cover of $G$.
Theorem 3.6.
Let $\Phi$ be a computable graph invariant and let $t(k)$ denote the maximal vertex cover number of a $k$-vertex graph $H$ such that $\widehat{\Phi}(H)\neq 0$. Then #$\#\textnormal{{ParIndSub}}(\Phi)$ can be solved in time $f(k)\cdot n^{t(k)+1}$ for some computable function $f$.
The theorem is shown as a consequence of [11, Theorem 1.1] in Appendix A.
4 Subgraph Basis versus Fourier Basis
While previous works (see, e.g., [15, 16, 44]) analyse the alternating enumerator directly to obtain hardness results, we observe that there is a close connection to the Fourier coefficients of the natural Boolean function representing $\Phi$. This allows rely on a plethora of existing results (see, e.g., [39]) on the Fourier expansion of Boolean functions to obtain hardness results for #$\#\textnormal{{IndSub}}(\Phi)$ for various different types of graph invariants $\Phi$.
We start by representing a graph invariant as a Boolean function. Let us fix some $k\geq 1$. Let $\Phi$ be a $k$-vertex graph invariant. We associate the Boolean function $f_{\Phi}$ of arity $\binom{k}{2}$ defined as follows. For ease of notation, we write $K\coloneqq K_{k}$ for the complete graph on $k$ in the remainder of this section. We define the function
$f_{\Phi}\colon\{-1,1\}^{E(K)}\to{\mathbb{R}}$
via
$f_{\Phi}(\boldsymbol{x})\coloneqq\Phi(K[\boldsymbol{x}])$
where $K[\boldsymbol{x}]\coloneqq K[S_{\boldsymbol{x}}]$ and $S_{\boldsymbol{x}}=\{e\in E(K)\mid\boldsymbol{x}_{e}=-1\}$.
Consistent with previous notation, for every $S\subseteq E(K)$, we write $\widehat{f}_{\Phi}(S)$ to denote the Fourier coefficient of $f_{\Phi}$ on $S$, i.e.,
$f_{\Phi}(\boldsymbol{x})=\sum_{S\subseteq E(K)}\widehat{f}_{\Phi}(S)x^{S}.$
To connect the alternating enumerator to the Fourier coefficients, we focus on inclusion-wise maximal edge sets evaluating to a non-zero value. Indeed, to obtain hardness results via the results from Section 3, we aim to find $k$-vertex graphs $H$ such that $\widehat{\Phi}(H)\neq 0$ and $H$ has large treewidth. Since the treewidth cannot decrease when adding edges, it is reasonable to focus on graphs with an inclusion-wise maximal edge set satisfying $\widehat{\Phi}(H)\neq 0$.
Let $S\subseteq E(K)$. We say that $S$ is $\widehat{f}$-maximal if $\widehat{f}(S)\neq 0$ and $\widehat{f}(S^{\prime})=0$ for every $S\subsetneq S^{\prime}\subseteq E(K)$. Similarly, we say $S$ is $\widehat{\Phi}$-maximal if $\widehat{\Phi}(K[S])\neq 0$ and $\widehat{\Phi}(K[S^{\prime}])=0$ for every $S\subsetneq S^{\prime}\subseteq E(K)$.
Lemma 4.1.
Let $k\geq 1$ and let $\Phi$ be a graph invariant on $k$-vertex graphs. Also let $S\subseteq E(K)$. Then $S$ is $\widehat{\Phi}$-maximal if and only if $S$ is $\widehat{f}_{\Phi}$-maximal.
Remark 4.2.
We stress that Lemma 4.1 does not hold if $S$ is not $\widehat{\Phi}$-maximal (or equivalently, not $\widehat{f}_{\Phi}$-maximal). For example, consider the property $\Psi$ which only contains the complete graph $K$, i.e., $\Psi(G)=1$ if and only if $G$ is isomorphic to $K$. Then $\widehat{\Psi}(H)\neq 0$ if and only if $H$ is isomorphic to $K$. On the other hand, $\operatorname{supp}(\widehat{f}_{\Psi})$ contains all subsets of $E(K)$ by the uncertainty principle (Lemma 2.3). Observe that $E(K)$ is the unique $\widehat{\Psi}$-maximal set and it is the unique $\widehat{f}_{\Psi}$-maximal set. A second example is also discussed in Remark 6.3.
For the proof of Lemma 4.1, we introduce another multilinear polynomial in ${\mathbb{R}}[(x_{e})_{e\in E(K)}]$ related to the property $\Phi$. Let us write $\boldsymbol{x}=(x_{e})_{e\in E(K)}$. We define the polynomial
$q_{\Phi}(\boldsymbol{x})\coloneqq\sum_{E\subseteq E(K)}\Phi(K[E])\cdot\prod_{e\in E}x_{e}\prod_{e\in E(K)\setminus E}(1-x_{e}).$ (9)
Let $e_{1},\ldots,e_{m}$ be an enumeration of $E(K)$. Note that $q_{\Phi}(a_{1},\ldots,a_{m})=\Phi(K[E])$ if $a_{1},\ldots,a_{m}$ are the zero-one values defined by $a_{i}=[e_{i}\in E]$ for all $i\in[m]$.
Lemma 4.3.
Let $\Phi$ be a graph invariant on $k$-vertex graphs. Then
$q_{\Phi}(\boldsymbol{x})=\sum_{S\subseteq E(K)}\widehat{\Phi}(K[S])\cdot x^{S}.$ (10)
Proof.
For $E\subseteq E(K)$ let $q_{E}(\boldsymbol{x})\coloneqq\prod_{e\in E}x_{e}\prod_{e\in E(K)\setminus E}(1-x_{e})$. Then
$q_{E}(\boldsymbol{x})=\sum_{S\subseteq E(K)\setminus E}(-1)^{|S|}x^{E\cup S}=\sum_{S\supseteq E}(-1)^{|S\setminus E|}x^{S}.$
It follows that the coefficient of $x^{S}$ is equal to
$\sum_{E\subseteq S}(-1)^{|S\setminus E|}\cdot\Phi(K[E])=(-1)^{|S|}\cdot\sum_{E\subseteq S}(-1)^{E}\cdot\Phi(K[E])=\widehat{\Phi}(K[S]).$
Proof of Lemma 4.1.
Let us number the variables $\boldsymbol{x}=(x_{e})_{e\in E(K)}$ as $\boldsymbol{x}=(x_{1},\dots,x_{m})$ where $m\coloneqq|E(K)|=\binom{k}{2}$. We claim that
$q_{\Phi}(x_{1},\dots,x_{m})=f_{\Phi}(1-2x_{1},\dots,1-2x_{m}).$ (11)
First observe that, together with Equations (4) and (10), this implies the claim since the “inclusionwise maximal monomials” in both polynomials $q_{\Phi}(\boldsymbol{x})$ and (the Fourier expansion of) $f_{\Phi}(\boldsymbol{x})$ are identical.
To prove Equation (11), consider some $a_{1},\dots,a_{m}\in\{0,1\}$. Let $E\coloneqq\{i\in[m]\mid a_{i}=1\}$. Slightly abusing notation, we also view $E$ as a subset of $E(K)$ via the numbering of variables fixed at the beginning of the proof. Then
$q_{\Phi}(a_{1},\dots,a_{m})=\Phi(K[E])=f_{\Phi}(1-2a_{1},\dots,1-2a_{m})$
since the function $t(y)=1-2y$ satisfies $t(0)=1$ and $t(1)=-1$. Since $f_{\Phi}$ is the unique multilinear polynomial satisfying all these equations, we conclude Equation (11). ∎
5 Hardness Results via the Fourier Basis
With Lemma 4.1 at hand, we can now exploit tools from the analysis of Boolean functions to find graphs with non-zero alternating enumerator. Together with the known results discussed in Section 3, this allows us to obtain hardness results for the problems #$\#\textnormal{{IndSub}}(\Phi)$ and #$\#\textnormal{{ParIndSub}}(\Phi)$.
5.1 Moderately Small Properties
As a first application, we consider the uncertainty principle (see Lemma 2.3). For ease of notation, for a graph invariant $\Phi$ on $k$-vertex graphs, we write
$\operatorname{supp}(\Phi)\coloneqq\{K[S]\mid S\in\operatorname{supp}(f_{\Phi})\}.$
Observe that $\operatorname{supp}(\Phi)$ contains exactly the graphs $G$ with vertex set $[k]$ for which $\Phi(G)\neq 0$. For an arbitrary graph invariant $\Phi$ (which is not only defined on $k$-vertex graphs) we also write $\operatorname{supp}_{k}(\Phi)\coloneqq\operatorname{supp}(\Phi^{(k)})$.
Theorem 5.1.
For every $10<\varepsilon<$ 1 there is some $\delta>0$ such that the following holds. Let $k\geq 1$ and let $\Phi$ be a $k$-vertex graph invariant such that
$1\leq|\operatorname{supp}(\Phi)|\leq(2-\varepsilon)^{\binom{k}{2}}.$
Then there is a $k$-vertex graph $H$ such that $\widehat{\Phi}(H)\neq 0$ and $|E(H)|\geq\delta\cdot\binom{k}{2}$.
Proof.
Let $10<\varepsilon<$ 1. We define $\alpha\coloneqq\log_{2}(1+\frac{\varepsilon}{2})$. Note that $10<\alpha<$ 1. So there is a unique $0<\delta<\frac{1}{2}$ such that $\alpha=\mathbf{H}(\delta)$ where $\mathbf{H}$ is the entropy function (see Lemma 2.4).
Now let $k\geq 1$ and let $\Phi$ be a $k$-vertex graph invariant such that
$1\leq|\operatorname{supp}(\Phi)|\leq(2-\varepsilon)^{\binom{k}{2}}.$
Since $|\operatorname{supp}(\Phi)|\geq 1$ we get that $f_{\Phi}$ is not the zero function. By Lemma 2.3 it holds that
$|\operatorname{supp}(\widehat{f}_{\Phi})|$ $\geq 2^{\binom{k}{2}}\cdot|\operatorname{supp}(f_{\Phi})|^{-1}=2^{\binom{k}{2}}\cdot|\operatorname{supp}(\Phi)|^{-1}$
$\geq 2^{\binom{k}{2}}\cdot(2-\varepsilon)^{-\binom{k}{2}}>\left(1+\frac{\varepsilon}{2}\right)^{\binom{k}{2}}=2^{\alpha\cdot\binom{k}{2}}.$
So there is some $S\in\operatorname{supp}(\widehat{f}_{\Phi})$ such that $|S|\geq\delta\cdot\binom{k}{2}$ by Lemma 2.4. We may suppose without loss of generality that $S$ is an inclusion-wise maximal set from $\operatorname{supp}(\widehat{f}_{\Phi})$ since adding further elements only increases the cardinality. We set $H\coloneqq K_{k}[S]$. Then $\widehat{\Phi}(H)\neq 0$ by Lemma 4.1. ∎
By combining Theorem 5.1 with Theorems 3.4(a) and 3.5(a), we obtain the following.
Corollary 5.2 (Theorem 1.3 restated).
For every $10<\varepsilon<$ 1 there are $N_{0},\delta>0$ such that the following holds. Let $k\geq N_{0}$ and let $\Phi$ be a $k$-vertex graph invariant such that
$1\leq|\operatorname{supp}(\Phi)|\leq(2-\varepsilon)^{\binom{k}{2}}.$
Then there is no algorithm that solves #$\#\textnormal{{IndSub}}(\Phi)$ in time $O(n^{\delta\cdot k/\sqrt{\log k}})$ unless ETH fails.
Moreover, for every computable graph invariant $\Phi$ such that
$1\leq|\operatorname{supp}_{k}(\Phi)|\leq(2-\varepsilon)^{\binom{k}{2}}$
holds for infinitely many $k$, the problem #$\#\textnormal{{ParIndSub}}(\Phi)$ is #W[1]-hard.
A notable example of graph properties that have moderately small support are hereditary graph properties. A graph property $\Phi$ is hereditary if $G[X]\in\Phi$ for every graph $G$ such that $G\in\Phi$ and every $X\subseteq V(G)$, i.e., $\Phi$ is closed under taking (vertex-)induced subgraphs.
Theorem 5.3.
Let $\Phi$ be a non-trivial hereditary graph property. Then there is an integer $k_{0}\geq 1$ and some $\varepsilon>0$ such that
$|\operatorname{supp}_{k}(\Phi)|\leq(2-\varepsilon)^{\binom{k}{2}}$
for every $k\geq k_{0}$.
Proof.
Since $\Phi$ is non-trivial there is some graph $F$ such that $\Phi(F)=0$. So every graph $G\in\Phi$ does not contain $F$ as an induced subgraph. Now the statement follows directly from [40] that bounds the number of graphs that do not contain $F$ as an induced subgraph. ∎
So in particular, Corollary 5.2 generalizes the results from [20] giving #W[1]-hardness for non-trivial hereditary properties. However, we point out that [20] shows that #$\#\textnormal{{ParIndSub}}(\Phi)$ for hereditary properties $\Phi$ cannot be solved in time $n^{o(k)}$ assuming ETH whereas Corollary 5.2 only rules out algorithms with running time $n^{o(k/\sqrt{\log k})}$ (if $\Phi$ is not meager). On the other hand, we stress that Corollary 5.2 also covers many non-hereditary properties and even holds for graph invariants that are not necessarily 00-$1$-valued. In particular, negative weights are allowed, which could a priori lead to intricate cancellations.
5.2 Monotone Properties
Next, we concern ourselves with monotone properties. Let $k\geq 1$ and let $\Phi$ be a $k$-vertex graph property. We say that $\Phi$ is monotone if $G[S]\in\Phi$ for every $G\in\Phi$ and every $S\subseteq E(G)$. For an arbitrary property $\Phi$, we say that $\Phi$ is monotone on $k$-vertex graphs if $\Phi^{(k)}$ is monotone.
Remark 5.4.
In the literature (on the complexity of counting induced subgraphs; see, e.g., [16, 20, 45]) this type of property is usually called edge-monotone whereas monotone properties are usually those that are hereditary and edge-monotone (on $k$-vertex graphs for all $k\geq 1$).
We prefer monotone properties since we are generally aiming to follow the notation used for analysing Boolean functions (see, e.g., [39]).
The following theorem is a simple modification of a result from [14].
Theorem 5.5.
Let $k\geq 1$ and $p$ a prime number. Let $\Phi$ is a $k$-vertex graph property that is monotone and not $k$-trivial.
Then there is some $k$-vertex graph $H$ such that $\widehat{\Phi}(H)\neq 0\bmod p$ and the complete bipartite graph $K_{\ell,\ell}$ is a subgraph of $H$ for some $\ell>\frac{k}{p^{2}}$.
More concretely, building on earlier work of Rivest and Vuillemin [41], Dodis and Khanna [14] show that the polynomial $q_{\Phi}$ has degree $\Omega(k^{2})$ if $\Phi$ is a graph property on $k$-vertex graphs that is monotone and not $k$-trivial. More precisely, [14] shows that $q_{\Phi}$ has degree $\Omega(k^{2})$ when interpreted as a polynomial over the two-element field ${\mathbb{Z}}_{2}$. Analysing their arguments in more detail, it is in fact shown that $q_{\Phi}$ contains a monomial (with non-zero coefficient) that corresponds to a graph $H$ such that $K_{\ell,\ell}$ is a subgraph of $H$ for some $\ell>\frac{k}{4}$. A simple extension of the arguments gives the statement above for every prime $p\geq 3$.
Since [14] does not provide proof details and their arguments need to be slightly extended, we give a full proof here. Let us stress again that the following proof of Theorem 5.5 precisely follows the arguments from [14].
Toward the proof of Theorem 5.5, we require some additional tools. For a finite set $\Omega$ we write $\operatorname{Sym}(\Omega)$ to denote the symmetric group over ground set $\Omega$. For $m\geq 1$ we also write $S_{m}$ as a shortcut for of $\operatorname{Sym}([m])$.
Let $p$ be a prime number and $m\geq 1$. Let $f\colon\{0,1\}^{m}\to{\mathbb{Z}}_{p}$ be a function over ${\mathbb{Z}}_{p}$ of arity $m$. A permutation $\sigma\in S_{m}$ is an automorphism of $f$ if $f(\boldsymbol{a})=f(\boldsymbol{a}^{\sigma})$ for every $\boldsymbol{a}=(a_{1},\dots,a_{m})\in\{0,1\}^{m}$ where $\boldsymbol{a}^{\sigma}=(a_{\sigma(1)},\dots,a_{\sigma(m)})$ is the tuple obtained from $\boldsymbol{a}$ by permuting its entries according to $\sigma$. We define the automorphism group of $f$, denoted by $\operatorname{Aut}(f)$, to be the group of all automorphisms of $f$. We say that $f$ is transitive if $\operatorname{Aut}(f)$ is transitive, i.e., for every $i,j\in[m]$ there is some $\sigma\in\operatorname{Aut}(f)$ such that $\sigma(i)=j$.
Similar to before, the function $f$ can be written as a multilinear polynomial over ${\mathbb{Z}}_{p}$ as
$q(\boldsymbol{x})=q(x_{1},\dots,x_{m})\coloneqq\sum_{\boldsymbol{a}\in\{0,1\}^{m}}f(\boldsymbol{a})\cdot\prod_{i\colon a_{i}=1}x_{i}\prod_{i\colon a_{i}=0}(1-x_{i}).$
We write $\deg_{p}(f)$ to denote the degree of $q$ (over the field ${\mathbb{Z}}_{p}$).
Also, we write 0=($\vec{0}=(0,\dots,0)$ and 1=($\vec{1}=(1,\dots,1)$ to denote the all-zero and all-one vector of length $m$, respectively.
Lemma 5.6.
Let $f\colon\{0,1\}^{m}\to{\mathbb{Z}}_{p}$ be a function over ${\mathbb{Z}}_{p}$ where $p$ is prime and $m=p^{k}$ for some integer $k\geq 1$. Also suppose $f$ is transitive and $f(\vec{0})\neq f(\vec{1})$. Then $\deg_{p}(f)=m$.
Proof.
Let $q(\boldsymbol{x})=q(x_{1},\dots,x_{m})$ denote the unique multilinear polynomial over ${\mathbb{Z}}_{p}$ representing $f$. We write the polynomial as
$q(\boldsymbol{x})=\sum_{S\subseteq[m]}\widehat{q}(S)\cdot x^{S}$
for appropriate numbers $\widehat{q}(S)\in{\mathbb{Z}}_{p}$ for every $S\subseteq[m]$. Observe that $f(\vec{0})=q(\vec{0})=\widehat{q}(\emptyset)$ and
$f(\vec{1})=q(\vec{1})=\sum_{S\subseteq[m]}\widehat{q}(S).$
We need to argue that $\widehat{q}([m])\neq 0$. The following two claims form the key ingredients.
Claim 5.7.
For every $S\subseteq[m]$ and $\sigma\in\operatorname{Aut}(f)$, we have $\widehat{q}(S)=\widehat{q}(S^{\sigma})$, where $S^{\sigma}=\{\sigma(i)\mid i\in S\}$.
Proof.
Let $\sigma\in\operatorname{Aut}(f)$. Consider the polynomial
$q_{\sigma}(\boldsymbol{x})=\sum_{S\subseteq[m]}\widehat{q}(S^{\sigma})\cdot x^{S}.$
We have that
$q(\boldsymbol{a})=f(\boldsymbol{a})=f(\boldsymbol{a}^{\sigma})=q_{\sigma}(\boldsymbol{a})$
for all $\boldsymbol{a}\in\{0,1\}^{m}$. Since the multilinear representation of $f$ is unique, we conclude that $q=q_{\sigma}$ which implies that $\widehat{q}(S)=\widehat{q}(S^{\sigma})$ for every $S\subseteq[m]$. ∎
Claim 5.8.
For every $\emptyset\subsetneq S\subsetneq[m]$, the cardinality of $S^{\operatorname{Aut}(f)}\coloneqq\{S^{\sigma}\mid\sigma\in\operatorname{Aut}(f)\}$ is divisible by $p$.
Proof.
Let $\emptyset\subsetneq S\subsetneq[m]$. We show that
$|S|\cdot|S^{\operatorname{Aut}(f)}|=m\cdot|\{S^{\prime}\in S^{\operatorname{Aut}(f)}\mid 1\in S^{\prime}\}|.$ (12)
First observe that this implies the claim, since $|S|<m$ and hence, at least one prime factor of $m=p^{k}$ needs to divide $|S^{\operatorname{Aut}(f)}|$.
To show Equation (12), consider the $|S^{\operatorname{Aut}(f)}|$ by $m$ matrix $M$ with entries
$M_{S^{\prime},i}=\begin{cases}1&\text{if }i\in S^{\prime}\\ 0&\text{otherwise}\end{cases}$
for all $S^{\prime}\in S^{\operatorname{Aut}(f)}$ and $i\in[m]$. Now, we count the number of ones in the matrix $M$ by rows and by columns.
First, counting the number of ones by rows gives $|S|\cdot|S^{\operatorname{Aut}(f)}|$ since each row contains exactly $|S|$ many ones since all sets in $S^{\operatorname{Aut}(f)}$ have the same size. On the other hand, the first column of $M$ contains |$|\{S^{\prime}\in S^{\operatorname{Aut}(f)}\mid 1\in S^{\prime}\}|$ many ones. Since $f$ is transitive, all columns contain the same number of ones, so in total we get $m\cdot|\{S^{\prime}\in S^{\operatorname{Aut}(f)}\mid 1\in S^{\prime}\}|$ many one-entries. Together, this implies Equation (12). ∎
[…]
$\sum_{S\in O_{j}}\widehat{q}(S)=0$
since $p$ divides $|O_{j}|$ and all operations are carried out over ${\mathbb{Z}}_{p}$. Overall, this means that
$f(\vec{1})=q(\vec{1})=\sum_{S\subseteq[m]}\widehat{q}(S)=\widehat{q}(\emptyset)+\widehat{q}([m]).$
Because $f(\vec{0})=q(\vec{0})=\widehat{q}(\emptyset)$ and $f(\vec{0})\neq f(\vec{1})$, we conclude that $\widehat{q}([m])\neq 0$. ∎
Let $p$ be a prime number and $m\geq 1$. Let $f\colon\{0,1\}^{m}\to{\mathbb{Z}}_{p}$ be a function over ${\mathbb{Z}}_{p}$ of arity $m$. Consistent with previous notation, for $J\subseteq[m]$ and $\boldsymbol{a}=(a_{j})_{j\in\overline{J}}\in\{0,1\}^{\overline{J}}$, we write $f_{J|\boldsymbol{a}}$ to denote the restriction of $f$ to $J$ using $\boldsymbol{a}$.
Let $q(\boldsymbol{x})=q(x_{1},\dots,x_{m})$ denote the unique multilinear polynomial over ${\mathbb{Z}}_{p}$ representing $f$. We write the polynomial as
$q(\boldsymbol{x})=\sum_{S\subseteq[m]}\widehat{q}(S)\cdot x^{S}$
for appropriate numbers $\widehat{q}(S)\in{\mathbb{Z}}_{p}$ for every $S\subseteq[m]$. We also use $q_{J|\boldsymbol{a}}$ to denote the polynomial with variables $(x_{j})_{j\in J}$ obtained from restricting $q$ to $J$ using $\boldsymbol{a}$. Observe that $q_{J|\boldsymbol{a}}$ is the unique multilinear representation of $f_{J|\boldsymbol{a}}$.
Let $\boldsymbol{x}_{J}\coloneqq(x_{j})_{j\in J}$ and suppose we write
$q_{J|\boldsymbol{a}}(\boldsymbol{x}_{J})=\sum_{S\subseteq J}\widehat{q}_{J|\boldsymbol{a}}(S)\cdot x^{S}.$
Then
$\widehat{q}_{J|\boldsymbol{a}}(S)=\sum_{T\subseteq\overline{J}}\widehat{q}(S\cup T)\prod_{j\in T}a_{j}.$ (13)
for all $S\subseteq J$ (cf. Equation (5)).
Lemma 5.9.
Let $p$ be a prime number and suppose $k=p^{\ell}$ for some integer $\ell\geq 1$. Also suppose $\Phi$ is a graph property on $k$-vertex graphs that is monotone and not $k$-trivial.
Then there is some $k$-vertex graph $H$ such that $\widehat{\Phi}(H)\neq 0\bmod p$ and the complete bipartite graph $K_{p^{\ell-1},p^{\ell-1}}$ is a subgraph of $H$.
Proof.
For $i\in\{0,\dots,\ell\}$ we define an equivalence relation $\sim_{i}$ on $[k]$ via
$a\sim_{i}b\quad\iff\quad\left\lceil\frac{a}{p^{i}}\right\rceil=\left\lceil\frac{b}{p^{i}}\right\rceil.$
Also, we set $F_{i}$ to be the graph with vertex set $V(F_{i})\coloneqq[k]$ and edge set $E(F_{i})\coloneqq\{ab\mid a\neq b\wedge a\sim_{i}b\}$. Note that $F_{i}$ is a disjoint union of $p^{\ell-i}$ many complete graphs of size $p^{i}$. Moreover,
$\emptyset=E(F_{0})\subseteq E(F_{1})\subseteq\dots\subseteq E(F_{\ell})=\binom{[k]}{2}.$
Since $\Phi$ is monotone and not $k$-trivial, there is some $i^{*}\in\{0,\dots,\ell-1\}$ such that $\Phi(F_{i})=1$ for all $i\leq i^{*}$, and $\Phi(F_{i})=0$ for all $i>i^{*}$.
[…]
Figure 1: Visualization of the construction of the graph $H^{*}$ in the proof of Lemma 5.9 for $p=3$. Each thick gray edge represents a complete biclique between the corresponding sets.
We define $H_{0}\coloneqq F_{i^{*}}$ and define a second sequence of graphs (see also Figure 1). We partition the set $[k]$ into $p$ intervals $A_{1},\dots,A_{p}$ of size $p^{\ell-1}$ each, i.e.,
$A_{i}\coloneqq\left\{(i-1)\cdot p^{\ell-1}+1,\dots,i\cdot p^{\ell-1}\right\}.$
Also, fix an arbitrary order $\{e_{1},\dots,e_{m}\}=\{vw\mid v\neq w\in[p]\}=\binom{[p]}{2}$ on the two-elements subsets of $[p]$, and define $M_{j}\coloneqq\{e_{1},\dots,e_{j}\}$ for all $j\in[p]$. We define the graph $H_{j}$ with vertex set $V(H_{j})=[k]$ and edge set
$E(H_{j})\coloneqq E(H_{0})\cup\{ab\mid a\in A_{v},b\in A_{w},vw\in M_{j}\}.$
Observe that
$E(H_{0})\subseteq E(H_{1})\subseteq\dots\subseteq E(H_{p}).$
Also $\Phi(H_{0})=1$ and $\Phi(H_{p})=0$ since $H_{p}$ contains a subgraph isomorphic to $F_{i^{*}+1}$. So there is some $j^{*}\in\{0,\dots,p-1\}$ such that $\Phi(H_{j})=1$ for all $j\leq j^{*}$, and $\Phi(H_{j})=0$ for all $j>j^{*}$. Let us set $H^{*}\coloneqq H_{j^{*}}$ and $V\coloneqq A_{v}$ and $W\coloneqq A_{w}$ where $vw=e_{j^{*}}$.
Claim 5.10.
For every $i\in[p]$ and every $s,s^{\prime}\in A_{i}$ there is an automorphism $\sigma\in\operatorname{Aut}(H^{*})$ such that $\sigma(s)=s^{\prime}$ and $\sigma(t)=t$ for all $t\in[k]\setminus A_{i}$.
Proof.
This follows immediately from observing that $H^{*}[A_{i}]$ is a disjoint union of complete graphs of the same size, and every $s,s^{\prime}\in A_{i}$ have the same neighbors outside of $A_{i}$ in $H^{*}$. ∎
Now, consider the polynomial $q_{\Phi}(\boldsymbol{x})$ defined over ${\mathbb{Z}}_{p}$ where $\boldsymbol{x}=(x_{e})_{e\in E(K_{k})}$, and also let $f\colon\{0,1\}^{E(K)}\to{\mathbb{Z}}_{p}$ be the function represented by $q_{\Phi}$. Let
$J\coloneqq\{vw\in E(K_{k})\mid v\in V,w\in W\}$
and $\overline{J}\coloneqq E(K_{k})\setminus J$. Let $\boldsymbol{a}=(a_{e})_{e\in\overline{J}}\in\{0,1\}^{\overline{J}}$ be the tuple defined via
$a_{e}\coloneqq\begin{cases}1&\text{if }e\in E(H^{*})\\ 0&\text{otherwise}\end{cases}.$
Now, consider the polynomial $q^{*}=(q_{\Phi})_{J|\boldsymbol{a}}$ and the function $f^{*}\colon\{0,1\}^{J}\to{\mathbb{Z}}_{p}$ represented by $q^{*}$.
We claim that $f^{*}$ satisfies the requirements of Lemma 5.6. First, $|J|=|V\times W|=p^{\ell-1}\cdot p^{\ell-1}=p^{2\ell-2}$ is a prime power of $p$. Also, $f^{*}(\vec{0})=\Phi(H^{*})=1$ and $f^{*}(\vec{1})=\Phi(H_{j^{*}+1})=0$, so $f^{*}(\vec{0})\neq f^{*}(\vec{1})$. It remains to verify that $f^{*}$ is transitive.
Claim 5.11.
$f^{*}$ is transitive.
Proof.
Let $e_{1},e_{2}\in J$ and suppose $e_{1}=s_{1}t_{1}$ and $e_{2}=s_{2}t_{2}$ where $s_{1},s_{2}\in V$ and $t_{1},t_{2}\in W$. By Claim 5.10 there is automorphism $\sigma\in\operatorname{Aut}(H^{*})$ such that $\sigma(s_{1})=s_{2}$ and $\sigma(t_{1})=t_{2}$, and additionally $V=V^{\sigma}\coloneqq\{\sigma(s)\mid s\in V\}$ and $W=W^{\sigma}\coloneqq\{\sigma(t)\mid t\in W\}$.
Now let $\bar{\sigma}\in\operatorname{Sym}(E(K_{k}))$ be defined via $\bar{\sigma}(st)\coloneqq\sigma(s)\sigma(t)$. Observe that $\bar{\sigma}(e_{1})=e_{2}$ and $J^{\bar{\sigma}}=J$. We claim that $\bar{\sigma}|_{J}\in\operatorname{Aut}(f^{*})$. Indeed, $\bar{\sigma}\in\operatorname{Aut}(f)$ since $f$ represents a graph property. Also, $\boldsymbol{a}^{\bar{\sigma}}=(a_{\sigma(e)})_{e\in\overline{J}}=\boldsymbol{a}$ since $\sigma\in\operatorname{Aut}(H^{*})$. Together, this implies that $\bar{\sigma}|_{J}\in\operatorname{Aut}(f^{*})$. ∎
By Lemma 5.6 we get that $\deg_{p}(f^{*})=|J|$. By Equation (13) and Lemma 4.3 this implies that there is some $J\subseteq S\subseteq E(K_{k})$ such that $\widehat{\Phi}(K_{k}[S])\neq 0\bmod p$. Hence, we can set $H\coloneqq K_{k}[S]$. Observe that $|V|=|W|=k/p=p^{\ell-1}$ and the complete bipartite graph between $V$ and $W$ is a subgraph of $H$ since $J\subseteq E(H)$. ∎
Proof of Theorem 5.5.
Let $\ell$ be maximal such that $p^{\ell}\leq k$. We prove by induction on $d=k-p^{\ell}$ that there is some $k$-vertex graph $H$ such that $\widehat{\Phi}(H)\neq 0\bmod p$ and $K_{p^{\ell-1},p^{\ell-1}}$ is a subgraph of $H$.
For the base case $d=0$ the statement immediately follows from Lemma 5.9. So suppose $d>0$. We distinguish three cases.
  • First suppose that $\Phi(K_{1,k-1})=1$. We define the property $\Psi$ on $(k-1)$-vertex graphs via $\Psi(G)=1$ if and only if $\Phi(G^{+})=1$ where $G^{+}$ is the graph obtained from $G$ by adding a universal vertex. Then $\Psi$ is a graph property that is monotone and non-trivial on $(k-1)$-vertex graphs. Indeed, observe that $\Psi(I_{k-1})=\Phi(K_{1,k-1})=1$, where $I_{k-1}$ is the edge-less graph on $k-1$ vertices, and $\Psi(K_{k-1})=\Phi(K_{k})=0$.
    By the induction hypothesis, we conclude that there is some $(k-1)$-vertex graph $H^{\prime}$ such that $\widehat{\Psi}(H^{\prime})\neq 0\bmod p$ and $K_{p^{\ell-1},p^{\ell-1}}$ is a subgraph of $H^{\prime}$.
    Now, let $J\coloneqq\{e\in E(K_{k})\mid k\notin e\}$. Then
    $q_{\Psi}=(q_{\Phi})|_{J|\vec{1}}.$
    Using Equation (13) and Lemma 4.3, we conclude that there is some $k$-vertex graph $H$ such that $\widehat{\Phi}(H)\neq 0\bmod p$ and $H^{\prime}\subseteq H$.
  • Next suppose that $\Phi(K_{k-1}+K_{1})=0$ where $K_{k-1}+K_{1}$ is the disjoint union of the complete graphs $K_{k-1}$ and $K_{1}$. We define the property $\Psi$ via $\Psi(G)=1$ if and only if $\Phi(G+K_{1})=1$ where $G+K_{1}$ is the graph obtained from $G$ by adding an isolated vertex. Then $\Psi$ is a graph property that is monotone and non-trivial on $(k-1)$-vertex graphs. Indeed, observe that $\Psi(I_{k-1})=\Phi(I_{k})=1$ and $\Psi(K_{k-1})=\Phi(K_{k-1}+K_{1})=0$.
    By the induction hypothesis, we conclude that there is some $(k-1)$-vertex graph $H^{\prime}$ such that $\widehat{\Psi}(H^{\prime})\neq 0\bmod p$ and $K_{p^{\ell-1},p^{\ell-1}}$ is a subgraph of $H^{\prime}$.
    Now, let $J\coloneqq\{e\in E(K_{k})\mid k\notin e\}$. Then
    $q_{\Psi}=(q_{\Phi})|_{J|\vec{0}}.$
    Using Equation (13) and Lemma 4.3, we conclude that there is some $k$-vertex graph $H$ such that $\widehat{\Phi}(H)\neq 0\bmod p$ and $H^{\prime}\subseteq H$.
  • Finally, suppose neither of the previous two cases holds. Let $r\coloneqq k-2p^{\ell-1}$. Observe that $r>0$ since $2p^{\ell-1}<p^{\ell}$. Let $s\coloneqq p^{\ell-1}$. Consider the graph $H^{*}=K_{r+s}+I_{s}$ on vertex set $[k]$, i.e., $H^{*}$ is a complete graph on $r+s$ vertices together with $s$ isolated vertices. Let $U,V,W$ be a partition of $[k]$ so that $|U|=r$, $|V|=|W|=s$ and $H^{*}[U\cup V]$ is a complete graph.
    Now, we proceed similarly to the proof of Lemma 5.9. Consider the polynomial $q_{\Phi}(\boldsymbol{x})$ defined over ${\mathbb{Z}}_{p}$ where $\boldsymbol{x}=(x_{e})_{e\in E(K_{k})}$. Let
    $J\coloneqq\{vw\in E(K_{k})\mid v\in V,w\in W\}$
    and $\overline{J}\coloneqq E(K_{k})\setminus J$. Let $\boldsymbol{a}=(a_{e})_{e\in\overline{J}}\in\{0,1\}^{\overline{J}}$ be the tuple defined via
    $a_{e}\coloneqq\begin{cases}1&\text{if }e\in E(H^{*})\\ 0&\text{otherwise}\end{cases}.$
    Now, consider the polynomial $q^{*}=(q_{\Phi})_{J|\boldsymbol{a}}$ and the function $f^{*}\colon\{0,1\}^{J}\to{\mathbb{Z}}_{p}$ represented by $q^{*}$.
    We claim that $f^{*}$ satisfies the requirements of Lemma 5.6. First, $|J|=p^{\ell-1}\cdot p^{\ell-1}=p^{2\ell-2}$ is a prime power of $p$. Also, $f^{*}(\vec{0})=1$ since otherwise $\Phi(K_{1,k-1})=1$ by monotonicity which leads to the first case. Moreover, $f^{*}(\vec{1})=0$ since otherwise $\Phi(K_{k-1}+K_{1})=0$ by monotonicity which leads to the second case. In particular, $f^{*}(\vec{0})\neq f^{*}(\vec{1})$. Also, $f^{*}$ is transitive by the same arguments as in Claim 5.11
    By Lemma 5.6 we get that $\deg_{p}(f^{*})=|J|$. By Equation (13) and Lemma 4.3 this implies that there is some $J\subseteq S\subseteq E(K_{k})$ such that $\widehat{\Phi}(K_{k}[S])\neq 0\bmod p$. Hence, we can set $H\coloneqq K_{k}[S]$. Observe that $|V|=|W|=s=p^{\ell-1}$ and the complete bipartite graph between $V$ and $W$ is a subgraph of $H$ since $J\subseteq E(H)$.
This completes the induction. To finish the proof we note that $p^{\ell-1}>\frac{k}{p^{2}}$. ∎
By combining Theorem 5.5 for $p=2$ with Theorems 3.4(a) and 3.5(b), we obtain the following corollary which improves on the results obtained in [16].
Corollary 5.12.
There are $N_{0},\delta>0$ such that the following holds. Let $k\geq N_{0}$ and let $\Phi$ be a $k$-vertex graph property that is monotone and not $k$-trivial. Then there is no algorithm that solves #$\#\textnormal{{IndSub}}(\Phi)$ in time $O(n^{\delta\cdot k})$ unless ETH fails.
Moreover, for every computable graph property $\Phi$ that is monotone on $k$-vertex graphs and not $k$-trivial for infinitely many $k$, the problem #$\#\textnormal{{ParIndSub}}(\Phi)$ is #W[1]-hard.
Also, by combining Theorem 5.5 with Theorems 3.4(b) and 3.5(c), we obtain hardness for modular counting.
Corollary 5.13 (Theorem 1.4 restated).
Let $p$ be a fixed prime. Then there are $N_{0},\delta>0$ such that the following holds. Let $k\geq N_{0}$ and let $\Phi$ be a $k$-vertex graph property that is monotone and not $k$-trivial. Then there is no algorithm that solves #$\#_{p}\textnormal{{IndSub}}(\Phi)$ in time $O(n^{\delta\cdot k})$ unless rETH fails.
[…]
5.3 Fully Symmetric Properties
We say that a $k$-vertex graph property $\Phi$ is fully symmetric if $\Phi(G)=\Phi(G^{\prime})$ for all $k$-vertex graphs $G,G^{\prime}$ with $|E(G)|=|E(G^{\prime})|$. In other words, $\Phi$ is fully symmetric if it only depends on the number of edges. As usual, we say a graph property $\Phi$ is fully symmetric on $k$-vertex graphs if $\Phi^{(k)}$ is fully symmetric. After a series of partial results on the hardness of fully symmetric properties [26, 28, 44], it was shown in [45] that #$\#\textnormal{{IndSub}}(\Phi)$ is hard for every non-trivial, fully symmetric $k$-vertex graph property. However, [45] could only rule out algorithms with running time $n^{o(k/\sqrt{\log k})}$ for #$\#\textnormal{{IndSub}}(\Phi)$ assuming ETH. In the following, we give a simpler proof of hardness and obtain lower bounds that are tight under ETH.
Our proof relies on the following theorem from [48]. A Boolean function $f\colon\{-1,1\}^{m}\to{\mathbb{R}}$ is fully symmetric if $f(\boldsymbol{a})=f(\boldsymbol{a}^{\sigma})$ for every $\boldsymbol{a}\in\{-1,1\}^{m}$ and every $\sigma\in S_{m}$, i.e., $f(\boldsymbol{a})$ only depends on the number of $1$-entries in $\boldsymbol{a}$. Observe that a property $\Phi$ on $k$-vertex graphs is fully symmetric if and only if $f_{\Phi}$ is fully symmetric.
Theorem 5.14 ([48, Theorem 2.8]).
Let $f\colon\{-1,1\}^{m}\to\{0,1\}$ be a Boolean function that is fully symmetric and non-constant. Then $\deg(f)\geq p-1$ where $p$ is the largest prime number such that $p\leq m+1$.
We note that [48, Theorem 2.8] is stated in a slightly different form, but the version stated here immediately follows from the proof. We also remark that for every $m\geq 2$ there is a prime $p\leq m$ such that $p\geq m-O(m^{0.525})$ using known estimates on the maximal gap between consecutive prime numbers [5].
Theorem 5.15.
Let $k\geq 2$. Suppose $\Phi$ is a $k$-vertex graph property that is fully symmetric and not $k$-trivial. Then there is a $k$-vertex graph $H$ such that $\widehat{\Phi}(H)\neq 0$ and $K_{\ell}$ is a subgraph of $H$ where $\ell\geq 2$ is maximal such that
$\binom{\ell}{2}+1\leq p\leq\binom{k}{2}+1$
for some prime $p$.
Proof.
Let $m\coloneqq p-1$ and note that $m\geq 1$. Then
$\binom{\ell}{2}\leq m\leq\binom{k}{2}.$
So there is a set $J\subseteq E(K_{k})$ such that $|J|=m$ and $E(K_{\ell})\subseteq J$. Let us write $\overline{J}\coloneqq E(K_{k})\setminus J$.
Claim 5.16.
There is a tuple $\boldsymbol{z}\in\{-1,1\}^{\overline{J}}$ such that the restriction
$f^{*}=(f_{\Phi})_{J|\boldsymbol{z}}\colon\{-1,1\}^{J}\to\{0,1\}$
of $f_{\Phi}$ to $J$ using $\boldsymbol{z}$ is fully symmetric and non-constant.
Proof.
We first show that $f^{*}=(f_{\Phi})_{J|\boldsymbol{z}}$ is fully symmetric for every tuple $\boldsymbol{z}\in\{-1,1\}^{\overline{J}}$. Indeed, let $\sigma\in\operatorname{Sym}(J)$ be a permutation of $J$, and let $\sigma^{*}\in\operatorname{Sym}(E(K_{k}))$ be the unique extension of $\sigma$ that fixes all points outside of $J$. Then $f^{*}(\boldsymbol{a})=f_{\Phi}(\boldsymbol{a},\boldsymbol{z})=f_{\Phi}((\boldsymbol{a},\boldsymbol{z})^{\sigma^{*}})=f_{\Phi}((\boldsymbol{a}^{\sigma},\boldsymbol{z}))=f^{*}(\boldsymbol{a}^{\sigma})$ where the second equality holds since $f_{\Phi}$ is fully symmetric.
So it remains to argue that there is some $\boldsymbol{z}\in\{-1,1\}^{\overline{J}}$ such that $(f_{\Phi})_{J|\boldsymbol{z}}$ is non-constant. Suppose towards a contradiction that $(f_{\Phi})_{J|\boldsymbol{z}}$ is constant for every $\boldsymbol{z}\in\{-1,1\}^{\overline{J}}$. This means that for every $\boldsymbol{z}\in\{-1,1\}^{\overline{J}}$ there is some $\alpha(\boldsymbol{z})\in\{0,1\}$ such that $(f_{\Phi})_{J|\boldsymbol{z}}$ always evaluates to $\alpha(\boldsymbol{z})$. Since $f_{\Phi}$ is non-constant, there are $\boldsymbol{z}_{1},\boldsymbol{z}_{2}\in\{-1,1\}^{\overline{J}}$ such that $\alpha(\boldsymbol{z}_{1})\neq\alpha(\boldsymbol{z}_{2})$. Additionally, we may assume that $\boldsymbol{z}_{1}$ and $\boldsymbol{z}_{2}$ only differ in one entry. But now there are $\boldsymbol{y}_{1},\boldsymbol{y}_{2}\in\{-1,1\}^{J}$ such that $(\boldsymbol{y}_{1},\boldsymbol{z}_{1})$ and $(\boldsymbol{y}_{2},\boldsymbol{z}_{2})$ have the same number of $1$-entries since $m\geq 1$. So $f_{\Phi}(\boldsymbol{y}_{1},\boldsymbol{z}_{1})=\alpha(\boldsymbol{z}_{1})\neq\alpha(\boldsymbol{z}_{2})=f_{\Phi}(\boldsymbol{y}_{2},\boldsymbol{z}_{2})$. But this is a contradiction since $f_{\Phi}$ is fully symmetric. ∎
So $\deg(f^{*})=m$ by Theorem 5.14 which implies that $\widehat{f^{*}}(J)\neq 0$. By Equation (5) there is some set $T\subseteq\overline{J}$ such that $\widehat{f}_{\Phi}(J\cup S)\neq 0$. We may suppose that $S$ is chosen to be inclusion-wise maximal. Then $\widehat{\Phi}(K_{k}[J\cup S])\neq 0$ by Lemma 4.1. To finish the proof observe that $K_{\ell}$ is a subgraph of $H\coloneqq K_{k}[J\cup S]$ since $E(K_{\ell})\subseteq J$. ∎
Note that, by Bertrand’s Postulate, it is always possible to choose $\ell\geq\frac{k}{2}$. In fact, using the estimate on the maximal gap between consecutive prime numbers [5] discussed above, it is possible to choose $\ell\geq k-O(k^{0.025})$. So by combining Theorem 5.15 with Theorems 3.4(a) and 3.5(b), we obtain the following.
Corollary 5.17 (Theorem 1.5 restated).
There are $N_{0},\delta>0$ such that the following holds. Let $k\geq N_{0}$ and let $\Phi$ be a $k$-vertex graph property that is fully symmetric and not $k$-trivial. Then there is no algorithm that solves #$\#\textnormal{{IndSub}}(\Phi)$ in time $O(n^{\delta\cdot k})$ unless ETH fails.
Moreover, for every computable graph property $\Phi$ that is fully symmetric on $k$-vertex graphs and not $k$-trivial for infinitely many $k$, the problem #$\#\textnormal{{ParIndSub}}(\Phi)$ is #W[1]-hard.
We note that the second part of the corollary has already been obtained in [45] using different arguments, but the first part has not been known prior to this work.
5.4 Invariants with Few Distinct Hamming Weights
In this section, we show that the polynomial interpretation of properties underlying the Fourier transform also entails lower bounds for properties (and even graph invariants) $\Phi$ that avoid a superlinear number of Hamming weights. In particular, our theorem allows us to recover [26, Theorem 3.4], which yields hardness when $\Phi$ holds only for $o(k^{2})$ distinct Hamming weights, and [45, Main Theorem 5], which applies whenever $\omega(k)$ Hamming weights are avoided. Our result even holds for graph invariants rather than properties, and it also rules out $n^{o(\beta(k)/\sqrt{\log\beta(k)})}$ time algorithms for #$\#\textnormal{{ParIndSub}}(\Phi)$ under ETH; see the defintion of $\beta$ below.
Given a $k$-vertex graph invariant $\Phi$, define its Hamming weight set as ${\sf hw}(\Phi)\coloneqq\{|E(H)|\mid H\in\Phi\}.$ In the following, and throughout this section, let $m\coloneqq\binom{k}{2}$. Then ${\sf hw}(\Phi)\subseteq\{0,\ldots,m\}$ is the set of all edge-counts (i.e, Hamming weights, when viewed as bitstrings) among the graphs in $\Phi$. We then define a value that counts the avoided Hamming weights, except when $\Phi$ is empty, as
$\beta_{\Phi}=\begin{cases}0&\text{if }|{\sf hw}(\Phi)|=0\\ \binom{k}{2}-|{\sf hw}(\Phi)|&\text{otherwise}.\end{cases}$
For a general graph invariant $\Phi$ and $k\in{\mathbb{N}}$, we define $\beta_{\Phi}(k)\coloneqq\beta_{\Phi^{(k)}}$.
Lemma 5.18.
For every $k$-vertex graph invariant $\Phi$ with $\Phi(G)\geq 0$ for all graphs $G$, there is a $k$-vertex graph $H$ with $\widehat{\Phi}(H)\neq 0$ and $|E(H)|\geq\beta_{\Phi}$.
Proof.
We assume that $|{\sf hw}(\Phi)|\notin\{0,m\}$, as the lemma holds vacuously otherwise. Let $q_{\Phi}(\boldsymbol{x})$ be the multivariate polynomial from Equation (9) representing $\Phi$ over $\{0,1\}^{m}$. If $q_{\Phi}(\boldsymbol{x})$ contains a monomial of degree $d$, then Lemma 4.3 yields $\widehat{\Phi}(H)\neq 0$ for some graph $H$ with $d$ edges. So it suffices to prove that $\deg(q_{\Phi})\geq m-|{\sf hw}(\Phi)|$.
We consider the symmetric polynomial
$q_{\Phi}^{\mathrm{symm}}(x_{1},\ldots,x_{m})=\sum_{\pi\in S_{m}}q_{\Phi}(x_{\pi(1)},\ldots,x_{\pi(m)}).$
For $a_{1},\ldots,a_{m}\in\{0,1\}$, we have $q_{\Phi}^{\mathrm{symm}}(a_{1},\ldots,a_{m})\neq 0$ if and only if $a_{1}+\ldots+a_{m}\in{\sf hw}(\Phi)$. Indeed, every $H$ with $\Phi(H)\neq 0$ and $|E(H)|=a_{1}+\ldots+a_{m}$ contributes a positive term to the sum on the right-hand side, and only such $H$ can contribute. Being positive, the values from different terms cannot cancel in the overall summation.
This symmetric polynomial can be “compressed” to a univariate one. More precisely, by [38, Lemma 3.2] (see also [35]), there exists a univariate polynomial $\tilde{q}$ such that, for all $a_{1},\ldots,a_{m}\in\{0,1\}$, we have
$\tilde{q}(a_{1}+\ldots+a_{m})=q_{\Phi}^{\mathrm{symm}}(a_{1},\ldots,a_{m}).$
For integers $z\in\{0,\ldots,m\}$, we thus have $\tilde{q}(z)\neq 0$ if and only if $z\in{\sf hw}(\Phi)$. Since $|{\sf hw}(\Phi)|\notin\{0,m\}$, we have that $\tilde{q}$ is not the zero polynomial. But since $\tilde{q}$ has $m-|{\sf hw}(\Phi)|$ roots, its degree is at least $m-|{\sf hw}(\Phi)|$. It follows that
$\deg(q_{\Phi})\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \deg(q_{\Phi}^{\mathrm{symm}})\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \deg(\tilde{q})\leavevmode\nobreak\ \geq\leavevmode\nobreak\ m-|{\sf hw}(\Phi)|\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \beta_{\Phi}.$
This proves the lemma via Lemma 4.3. ∎
In the following, we that the assumption of nonnegativity on the graph invariant $\Phi$ can be dropped by passing from $\Phi$ to its pointwise square and observing that this at most doubles the degree of the representing polynomial. More formally, for two graph invariants $\Phi_{1},\Phi_{2}$, the point-wise product is the graph invariant $\Psi=\Phi_{1}\cdot\Phi_{2}$ defined via $\Psi(G)=\Phi_{1}(G)\cdot\Phi_{2}(G)$.
Lemma 5.19.
Let $\Phi_{1},\Phi_{2}$ be graph invariants and let $\Psi\coloneqq\Phi_{1}\cdot\Phi_{2}$. Then, for every graph $H$,
$\widehat{\Psi}(H)=\sum_{\begin{subarray}{c}S_{1},S_{2}\subseteq E(H)\colon\\ S_{1}\cup S_{2}=E(H)\end{subarray}}\widehat{\Phi}_{1}(H[S_{1}])\cdot\widehat{\Phi}_{2}(H[S_{2}]).$
Proof.
Suppose $k\coloneqq|V(H)|$. Let $m\coloneqq\binom{k}{2}$ and let $q(\boldsymbol{x})\coloneqq q_{\Psi}(\boldsymbol{x})$ be the multilinear polynomial from (9) representing $\Psi$ over $\{0,1\}^{m}$. Likewise, let $q_{1}(\boldsymbol{x})\coloneqq q_{\Phi_{1}}(\boldsymbol{x})$ and $q_{2}(\boldsymbol{x})\coloneqq q_{\Phi_{2}}(\boldsymbol{x})$ be the multilinear polynomials representing $\Phi_{1}$ and $\Phi_{2}$.
Given a polynomial $r(\boldsymbol{x})$ with maximum degree $2$ per variable, let $r_{\textsf{red}}$ denote the polynomial obtained by replacing $x_{i}^{2}$ with $x_{i}$ in every monomial. We clearly have $r(\boldsymbol{a})=r_{\textsf{red}}(\boldsymbol{a})$ for all $\boldsymbol{a}\in\{0,1\}^{m}$. Since every multilinear polynomial in $m$ indeterminates is determined by its evaluations on all values $\boldsymbol{a}\in\{0,1\}^{m}$, we obtain $q(\boldsymbol{x})=(q_{1}(\boldsymbol{x})\cdot q_{2}(\boldsymbol{x}))_{\textsf{red}}$, since $q(\boldsymbol{a})=q_{1}(\boldsymbol{a})\cdot q_{2}(\boldsymbol{a})$ for all $\boldsymbol{a}\in\{0,1\}^{m}$. We now consider the coefficients of $q(\boldsymbol{x})$.
In the product $q_{1}(\boldsymbol{x})\cdot q_{2}(\boldsymbol{x})$, monomials $x^{S}$ represent multisets $S$ of edges from $E(H)$ with multiplicity at most $2$ per edge. Let $\widehat{q}_{i}(S)$ denote the coefficient of $x^{S}$ in $q_{i}(\boldsymbol{x})$ for $i\in\{1,2\}$; we have $\widehat{q}_{i}(S)=\widehat{\Phi}_{i}(K_{k}[S])$ by Lemma 4.3. By collecting monomials after polynomial multiplication, the coefficient of $x^{S}$ in $q_{1}(\boldsymbol{x})\cdot q_{2}(\boldsymbol{x})$ equals
$\sum_{\begin{subarray}{c}S_{1},S_{2}\subseteq S\colon\\ S_{1}\cup S_{2}=S\end{subarray}}\widehat{\Phi}_{1}(K_{k}[S_{1}])\cdot\widehat{\Phi}_{2}(K_{k}[S_{2}]),$ (14)
where the union $S_{1}\cup S_{2}$ is interpreted as a multiset union, i.e., counting copies. Collecting monomials after substitution, it follows that the coefficient $\widehat{q}(S)=\widehat{\Psi}(H[S])$ of $x^{S}$ in $q(\boldsymbol{x})=(q_{1}(\boldsymbol{x})\cdot q_{2}(\boldsymbol{x}))_{\textsf{red}}$ for a set (rather than multiset) $S\subseteq E(H)$ also equals (14), but with $S_{1}\cup S_{2}$ interpreted as a set union, i.e., not counting copies. This proves the lemma, as $\widehat{q}_{i}(S)=\widehat{\Phi}_{i}(K_{k}[S])$ and $\widehat{q}(S)=\widehat{\Psi}(K_{k}[S])$. ∎
Applying Lemma 5.19 to $\Psi\coloneqq\Phi\cdot\Phi$, we obtain the following variant of Lemma 5.18.
Theorem 5.20.
For every $k$-vertex graph invariant $\Phi$, there is a $k$-vertex graph $H$ with $\widehat{\Phi}(H)\neq 0$ and $|E(H)|\geq\beta_{\Phi}/2$.
Proof.
Let $m=\binom{k}{2}$. For $\Psi:=\Phi\cdot\Phi$, we have $\Psi(G)\geq 0$ for all graphs $G$. Moreover, the graph invariants $\Psi$ and $\Phi$ satisfy ${\sf hw}(\Psi)={\sf hw}(\Phi)$ and thus $\beta_{\phi}=\beta_{\Psi}$. By Lemma 5.18, there is a $k$-vertex graph $H^{\prime}$ with $\widehat{\Psi}(H^{\prime})\neq 0$ and $|E(H^{\prime})|\geq\beta_{\Psi}=\beta_{\Phi}$. By Lemma 5.19, the set $E(H^{\prime})$ can be written as $E(H^{\prime})=S_{1}\cup S_{2}$ with $\widehat{\Phi}(H^{\prime}[S_{1}])\neq 0$ and $\widehat{\Phi}(H^{\prime}[S_{2}])\neq 0$. It follows that one of $S_{1}$ or $S_{2}$ contains at least $\beta_{\Phi}/2$ edges; choose $H$ to be the corresponding graph $H^{\prime}[S_{1}]$ or $H^{\prime}[S_{2}]$. ∎
By combining Theorem 5.20 with Theorems 3.4(a) and 3.5(a), we obtain the following. For functions $f,g:{\mathbb{N}}\to{\mathbb{Q}}$, we say that $f\in\omega_{\mathrm{inf}}(g)$ if there is an infinite sequence $n_{0},n_{1},n_{2},\ldots$ such that $f(n_{i})\in\omega(g(n_{i}))$. Note that $f(n_{i})$ and $g(n_{i})$ take $i$ as input. Recall that $\beta_{\Phi}(k)=\beta_{\Phi^{(k)}}$, where $\Phi^{(k)}$ is the $k$-th slice of $\Phi$.
Corollary 5.21 (Theorem 1.6 restated).
There are $N_{0},\delta>0$ such that the following holds. Let $k\geq N_{0}$ and $20<d\leq k/$ 2, let $\Phi$ be a $k$-vertex graph invariant, and let
$\beta_{\Phi}\geq dk.$
Then there is no algorithm that solves #$\#\textnormal{{IndSub}}(\Phi)$ in time $O(n^{\delta\cdot d/\sqrt{\log d}})$ unless ETH fails.
Moreover, for every computable graph invariant $\Phi$ such that
$\beta_{\Phi}(k)\in\omega_{\mathrm{inf}}(k),$
the problem #$\#\textnormal{{ParIndSub}}(\Phi)$ is #W[1]-hard.
6 All Degrees are Even
In this section we consider the property $\Phi_{\sf even}$ defined via
$\Phi_{\sf even}(G)=\begin{cases}1&\text{if }\deg_{G}(v)\equiv 0\bmod 2\text{ for all }v\in V(G)\\ 0&\text{otherwise,}\end{cases}$
i.e., $\Phi_{\sf even}$ contains exactly those graphs where all vertices have even degree. This property turns out to be interesting since $\widehat{\Phi}_{\sf even}(H)=0$ for most graphs $H$. While #$\#\textnormal{{ParIndSub}}(\Phi_{\sf even})$ is still #W[1]-hard, this property allows us to also derive non-trivial upper bounds on the complexity. Moreover, the fact that $\widehat{\Phi}_{\sf even}(H)=0$ for most graphs $H$ surprisingly allows us to obtain hardness of #$\#\textnormal{{ParIndSub}}(\Phi)$ for other properties $\Phi$.
The starting point for our analysis is the observation that each slice $\Phi_{\sf even}^{(k)}$ forms a vector space over the $2$-element field ${\mathbb{Z}}_{2}$. Indeed, $\Phi_{\sf even}^{(k)}$ contains exactly those graphs that are in the cycle space of $K_{k}$. This allows us to use basic tools from linear algebra to determine exactly which graphs $H$ satisfy $\widehat{\Phi}_{\sf even}(H)\neq 0$.
Theorem 6.1.
The graphs $H$ satisfying $\widehat{\Phi}_{\sf even}(H)\neq 0$ are precisely the bipartite graphs.
Proof.
Let $H$ be a bipartite graph. For every $S\subseteq E(H)$ with $\Phi_{\sf even}(H[S])=1$ we have $|S|$ even, so every term in
$\widehat{\Phi}_{\sf even}(H)=\sum_{S\subseteq E(H)}(-1)^{|S|}\cdot\Phi_{\sf even}(H[S])$
is non-negative. Since the term $\Phi_{\sf even}(H[\emptyset])=1$ is strictly positive, we get $\widehat{\Phi}_{\sf even}(H)\neq 0$.
Next, let $H$ be a non-bipartite graph and let $C^{*}\subseteq E(H)$ be the edges of an odd cycle in $H$. Consider the cycle space of $H$, i.e., the set of all subgraphs in which all vertices have even degree:
${\mathcal{C}}\coloneqq\{S\subseteq E(H)\mid H[S]\in\Phi_{\sf even}\}.$
As noted above, ${\mathcal{C}}$ can be viewed as a subspace of ${\mathbb{Z}}_{2}^{E(H)}$ where each $C\in{\mathcal{C}}$ is identified with its corresponding indicator vector $v_{C}\in{\mathbb{Z}}_{2}^{E(H)}$. Slightly abusing notation, we identify $C$ with $v_{C}$.
Claim 6.2.
There is a basis $\{C_{1},\dots,C_{d}\}\subseteq{\mathcal{C}}$ such that $|C_{i}|$ is odd for every $i\in[d]$.
Proof.
Clearly, there is a basis ${\mathcal{B}}$ of ${\mathcal{C}}$ such that $C^{*}\in{\mathcal{B}}$. But then
${\mathcal{B}}^{*}\coloneqq\{C\mid C\in{\mathcal{B}},|C|\text{ is odd}\}\cup\{C\oplus C^{*}\mid C\in{\mathcal{B}},|C|\text{ is even}\}$
is also a basis of ${\mathcal{C}}$, where $C\oplus C^{*}$ denotes the symmetric difference of $C$ and $C^{*}$ (i.e., the corresponding vectors over ${\mathbb{Z}}_{2}$ are added). Every element of ${\mathcal{B}}^{*}$ has odd cardinality, as desired. ∎
Now let $\{C_{1},\dots,C_{d}\}$ be the basis constructed in the claim. Since $|C_{i}|$ is odd for every $i\in[d]$, we get that
$|\bigoplus_{i\in M}C_{i}|\equiv|M|\mod 2$
for every $M\subseteq[d]$. It follows that
$\widehat{\Phi}_{\sf even}(H)$ $=\sum_{S\subseteq E(H)}(-1)^{|S|}\cdot\Phi_{\sf even}(H[S])$
$=\sum_{C\in{\mathcal{C}}}(-1)^{|C|}$
$=\sum_{M\subseteq[d]}(-1)^{|\bigoplus_{i\in M}C_{i}|}$
$=\sum_{M\subseteq[d]}(-1)^{|M|}$
$=0$
where the last equality follows from the Binomial Theorem. ∎
Remark 6.3.
A weaker version of Theorem 6.1 can also be obtained via the Fourier expansion. Let $k\geq 1$ and let $\Psi\coloneqq\Phi_{\sf even}^{(k)}$. The set $\operatorname{supp}(f_{\Psi})$ forms a subspace of ${\mathbb{Z}}_{2}^{m}$ for $m\coloneqq\binom{k}{2}$. Then $\operatorname{supp}(\widehat{f}_{\Psi})$ is the complement space of this space, see [39, Proposition 3.11]. Since $\operatorname{supp}(f_{\Psi})$ corresponds to the cycle space of $K_{k}$, we obtain that $\operatorname{supp}(\widehat{f}_{\Psi})$ corresponds to the cut space of $K_{k}$. It follows that the $\widehat{f}_{\Psi}$-maximal sets are precisely the edge sets of complete bipartite graphs on $k$ vertices. By Lemma 4.1, we can therefore conclude that
  1. 1.
    $\widehat{\Phi}_{\sf even}(H)\neq 0$ for all complete bipartite graphs $H$, and
  2. 2.
    if $\widehat{\Phi}_{\sf even}(H)\neq 0$ then $H$ is bipartite.
We remark that this weaker statement is sufficient to obtain all results presented in the remainder of this section. Our Theorem 6.1 shows that all bipartite graphs $H$ are in the support of $\widehat{\Phi}_{\sf even}(H)$, not just the complete bipartite graphs.
By combining Theorem 5.15 with Theorems 3.4(a) and 3.5(b), we obtain the following.
Corollary 6.4.
There are $N_{0},\delta>0$ such that for every $k\geq N_{0}$ there is no algorithm that solves #$\#\textnormal{{IndSub}}(\Phi_{\sf even}^{(k)})$ in time $O(n^{\delta\cdot k})$ unless ETH fails. Moreover, #$\#\textnormal{{ParIndSub}}(\Phi_{\sf even})$ is #W[1]-hard.
In contrast to the results in Section 5, Theorem 6.1 also allows us to obtain a non-trivial upper bound on the complexity of #$\#\textnormal{{ParIndSub}}(\Phi_{\sf even})$. This follows from Theorem 3.6, by observing that each bipartite graph with $k$ vertices has vertex cover number at most $\lfloor k/2\rfloor$.
Corollary 6.5.
The problem #$\#\textnormal{{ParIndSub}}(\Phi_{\sf even})$ can be solved in time $f(k)\cdot n^{\lfloor k/2\rfloor+1}$ for some computable function $f$.
Observe that the algorithm from the last corollary achieves roughly a quadratic speed-up over the brute-force algorithm, which runs in time $f(k)\cdot n^{k}$ for some computable function $f$. We note that we are not aware of any other non-meager property $\Phi$ that is “fundamentally different” from $\Phi_{\sf even}$ for which a similar (or better) running time can be achieved. [^4] If we complement the property $\Phi_{\sf even}$, or we complement every graph in $\Phi_{\sf even}$, then Theorem 6.1 still holds and the same runtime can be achieved. We do not view such variants of the property $\Phi_{\sf even}$ as “fundamentally different”.
Next, we use Theorem 6.1 to obtain hardness results for #$\#\textnormal{{IndSub}}(\Phi)$ for other graph invariants. Indeed, Theorem 6.1 shows that the property #$\#\textnormal{{IndSub}}(\Phi,k,$) is “relatively easy” in the sense that only few graphs have non-zero alternating enumerator. Surprisingly, this fact allows us to prove hardness results for other graph invariants by exploiting Lemma 5.19.
More precisely, to obtain hardness results for #$\#\textnormal{{IndSub}}(\Phi)$ for a $k$-vertex graph invariant $\Phi$, the key idea is to take a point-wise product $\Psi\coloneqq\Phi_{\sf even}\cdot\Phi$. If $\Psi(H^{*})\neq 0$ for some $k$-vertex graph $H^{*}$ that is “far away” from being bipartite, we can conclude that there is some $k$-vertex graph $H$ with many edges such that $\widehat{\Phi}(H)\neq 0$ using Lemma 5.19. This approach is demonstrated in the following corollary.
Corollary 6.6.
For every $k$-vertex graph property $\Phi$ whose intersection with $\Phi_{\sf even}$ is non-empty and contains only bipartite graphs, there is a $k$-vertex graph $H$ such that $\widehat{\Phi}(H)\neq 0$ and $K_{\ell}$ is a subgraph of $H$ for some $\ell\geq\frac{k}{2}$.
Proof.
Let $\Psi\coloneqq\Phi\cdot\Phi_{\sf even}$. Note that $\Psi$ is also a graph property that corresponds to the intersection of $\Phi$ and $\Phi_{\sf even}$. Then every $k$-vertex graph $H\in\Psi$ has an even number of edges. So
$\widehat{\Psi}(K_{k})=\sum_{S\subseteq E(K_{k})}(-1)^{|S|}\cdot\Psi(K_{k}[S])=\sum_{S\subseteq E(K_{k})}\Psi(K_{k}[S])$
denotes the number of $k$-vertex graphs in $\Psi$. Since there is at least one such graph, we conclude that $\widehat{\Psi}(K_{k})\neq 0$.
By Lemma 5.19 there are sets $S,S_{\sf even}\subseteq E(K_{k})$ such that $S\cup S_{\sf even}=E(K_{k})$ and $\widehat{\Phi}(K_{k}[S])=1$ and $\widehat{\Phi}_{\sf even}(K_{k}[S_{\sf even}])=1$. We set $H\coloneqq K_{k}[S]$. Since $\widehat{\Phi}_{\sf even}(K_{k}[S_{\sf even}])=1$, we conclude that $K_{k}[S_{\sf even}]$ is bipartite by Theorem 6.1. Let $V,W$ denote a bipartition of $K_{k}[S_{\sf even}]$. Then $\binom{V}{2}\subseteq S$ and $\binom{W}{2}\subseteq S$ since $S\cup S_{\sf even}=E(K_{k})$. In particular, $H$ contains a clique of size at least $\frac{k}{2}$. ∎
We can also combine a similar approach with the Uncertainty Principle (see Lemma 2.3) to obtain a variant of Theorem 5.1.
Theorem 6.7.
For every $10<\varepsilon<$ 1 there are $\delta,k_{0}>0$ such that the following holds. Let $k\geq k_{0}$ and let $\Phi$ be a $k$-vertex graph invariant such that
$1\leq|\operatorname{supp}(\Phi\cdot\Phi_{\sf even})|\leq\left(\sqrt{2}-\varepsilon\right)^{\binom{k}{2}}.$
Then there is a $k$-vertex graph $H$ such that $\widehat{\Phi}(H)\neq 0$ and $|E(H)|\geq\delta\cdot\binom{k}{2}$.
Observe that, in comparison with Theorem 5.1, we only require $\Phi\cdot\Phi_{\sf even}$ to have small support, but we need a stronger upper bound on its size.
Toward the proof of Theorem 6.7, we require the followings tool. Let $G$ be a graph. A set $S\subseteq E(G)$ is an edge bipartization if $G-S\coloneqq G[E(G)\setminus S]$ is bipartite. The edge bipartization number of $G$ is the minimum integer $k\geq 0$ such that there is an edge bipartization $S\subseteq E(G)$ of size $|S|=k$.
Lemma 6.8.
For every $10<\varepsilon<$ 1 there are $\delta,k_{0}>0$ such that the following holds. Let $k\geq k_{0}$ and let ${\mathcal{C}}$ be a set of graphs with vertex set $[k]$ such that
$|{\mathcal{C}}|\geq\left(\sqrt{2}+\varepsilon\right)^{\binom{k}{2}}.$
Then there is some $G\in{\mathcal{C}}$ with edge bipartization number at least $\delta\cdot\binom{k}{2}$.
Proof.
Let $10<\varepsilon<$ 1 and suppose $k_{0}$ is a sufficiently large integer (depending on $\varepsilon$). Also let
$\alpha\coloneqq\log_{2}\left(1+\frac{\varepsilon}{\sqrt{2}\cdot 3}\right).$
Then $10<\alpha<$ 1 and there is some $0<\delta<\frac{1}{2}$ such that $\alpha=\mathbf{H}(\delta)$ where $\mathbf{H}$ is the entropy function.
Now let $k\geq k_{0}$. Also let ${\mathcal{B}}$ denote the set of bipartite graphs with vertex set $[k]$. By [18, Lemma 7] it holds that
$|{\mathcal{B}}|\leq\left(\sqrt{2}+\frac{\varepsilon}{3}\right)^{\binom{k}{2}}$
since $k\geq k_{0}$ is sufficiently large. Next, let ${\mathcal{H}}$ denote the set of graphs $H$ with vertex set $[k]$ and $|E(H)|\leq\delta\cdot\binom{k}{2}$. Then, by Lemma 2.4, we get
$|{\mathcal{H}}|\leq 2^{\mathbf{H}(\delta)\cdot\binom{k}{2}}=\left(1+\frac{\varepsilon}{\sqrt{2}\cdot 3}\right)^{\binom{k}{2}}.$
We define ${\mathcal{B}}\otimes{\mathcal{H}}$ to be the set of graphs $G$ with vertex set $[k]$ such that $E(G)=E(B)\cup E(H)$ for some $B\in{\mathcal{B}}$ and $H\in{\mathcal{H}}$. Then every graph $G$ with vertex set $[k]$ and edge bipartization number at most $\delta\cdot\binom{k}{2}$ is contained in ${\mathcal{B}}\otimes{\mathcal{H}}$. Also,
$|{\mathcal{B}}\otimes{\mathcal{H}}|\leq|{\mathcal{B}}|\cdot|{\mathcal{H}}|$ $\leq\left(\sqrt{2}+\frac{\varepsilon}{3}\right)^{\binom{k}{2}}\cdot\left(1+\frac{\varepsilon}{\sqrt{2}\cdot 3}\right)^{\binom{k}{2}}$
$=\left(\sqrt{2}+\frac{\varepsilon}{3}+\frac{\varepsilon}{3}+\frac{\varepsilon^{2}}{\sqrt{2}\cdot 9}\right)^{\binom{k}{2}}<\left(\sqrt{2}+\varepsilon\right)^{\binom{k}{2}}.$
This implies that ${\mathcal{C}}$ contains some graph with edge bipartization number at least $\delta\cdot\binom{k}{2}$. ∎
Proof of Theorem 6.7.
Fix some $10<\varepsilon<$ 1 and choose $\delta,k_{0}>0$ according to Lemma 6.8.
Let $\Psi\coloneqq\Phi\cdot\Phi_{\sf even}$ and suppose
$1\leq|\operatorname{supp}(\Psi)|\leq\left(\sqrt{2}-\varepsilon\right)^{\binom{k}{2}}.$
Since $|\operatorname{supp}(\Psi)|\geq 1$ we get that $f_{\Psi}$ is not the zero function. By Lemma 2.3 it holds that
$|\operatorname{supp}(\widehat{f}_{\Psi})|\geq 2^{\binom{k}{2}}\cdot|\operatorname{supp}(f_{\Psi})|^{-1}=2^{\binom{k}{2}}\cdot|\operatorname{supp}(\Psi)|^{-1}\geq 2^{\binom{k}{2}}\cdot\left(\sqrt{2}-\varepsilon\right)^{-\binom{k}{2}}\geq\left(\sqrt{2}+\varepsilon\right)^{\binom{k}{2}}.$
By Lemma 6.8, there is some $S\in\operatorname{supp}(\widehat{f}_{\Psi})$ such that $K_{k}[S]$ has edge bipartization number at least $\delta\cdot\binom{k}{2}$. We may suppose without loss of generality that $S$ is an inclusion-wise maximal set from $\operatorname{supp}(\widehat{f}_{\Psi})$ since adding further elements only increases the edge bipartization number. We set $F\coloneqq K_{k}[S]$. Then $\widehat{\Psi}(F)\neq 0$ by Lemma 4.1. Also, $F$ has edge edge bipartization number at least $\delta\cdot\binom{k}{2}$.
Now, by Lemma 5.19 there are sets $B,S_{\sf even}\subseteq E(F)$ such that $B\cup S_{\sf even}=E(F)$ and $\widehat{\Phi}(F[B])\neq 0$ and $\widehat{\Phi}_{\sf even}(F[S_{\sf even}])\neq 0$. We set $H\coloneqq F[B]$. Since $\widehat{\Phi}_{\sf even}(F[S_{\sf even}])=1$, we conclude that $F[S_{\sf even}]$ is bipartite by Theorem 6.1. So $B$ is an edge bipartization of $F$ and hence, $|B|\geq\delta\cdot\binom{k}{2}$. ∎
By combining Theorem 6.7 with Theorems 3.4(a) and 3.5(a), we obtain the following.
Corollary 6.9 (Theorem 1.7 restated).
For every $10<\varepsilon<$ 1 there are $N_{0},\delta>0$ such that the following holds. Let $k\geq N_{0}$ and let $\Phi$ be a $k$-vertex graph invariant such that
$1\leq|\operatorname{supp}(\Phi\cdot\Phi_{\sf even})|\leq(\sqrt{2}-\varepsilon)^{\binom{k}{2}}.$
Then there is no algorithm that solves #$\#\textnormal{{IndSub}}(\Phi)$ in time $O(n^{\delta\cdot k/\sqrt{\log k}})$ unless ETH fails.
Moreover, for every computable graph invariant $\Phi$ such that
$1\leq|\operatorname{supp}_{k}(\Phi\cdot\Phi_{\sf even})|\leq(\sqrt{2}-\varepsilon)^{\binom{k}{2}}$
holds for infinitely many $k$, the problem #$\#\textnormal{{ParIndSub}}(\Phi)$ is #W[1]-hard.
7 Weisfeiler-Leman Dimension
Besides hardness results for induced subgraph counts, our techniques also allow us to bound or even precisely determine the Weisfeiler-Leman dimension of such graph parameters. The $k$-dimensional Weisfeiler-Leman algorithm ($k$-WL) is a standard heuristic for testing isomorphism of graphs, but it also has surprising connections to various other areas of computer science, including finite model theory and machine learning [23, 24, 30, 36]. We omit a detailed description of the WL algorithm here and refer the interested reader to [23, 24, 30]. Indeed, the following characterization shown by Dvorák [17] is sufficient for our purposes.
A homomorphism from a graph $F$ to a graph $G$ is a mapping $\varphi\colon V(F)\to V(G)$ such that $\varphi(v)\varphi(w)\in E(G)$ for every edge $vw\in E(F)$. We write #$\#\mathrm{Hom}(F\to G)$ to denote the number of homomorphisms from $F$ to $G$. Then two graphs $G,G^{\prime}$ are distinguished by $k$-WL if #$\#\mathrm{Hom}(F\to G)\neq\#\mathrm{Hom}(F\to G^{\prime})$ for some graph $F$ with $\operatorname{tw}(F)\leq k$.
The Weisfeiler-Leman dimension of a graph invariant $f$ is the minimal $k\geq 1$ such that $k$-WL distinguishes between all graphs $G$ and $G^{\prime}$ for which $f(G)\neq f(G^{\prime})$. For the graph parameters considered in this paper, such a number $k$ indeed exists. Due to the connections of WL to various other areas of computer science, the WL-dimension has become a widely studied measure (see, e.g., [3, 4, 21, 22, 32, 37]). Our results yield bounds on the WL-dimension of $f(\star)=\#\mathrm{IndSub}(\Phi^{(k)}\to\,\star\,)$ for various graph invariants $\Phi$.
More precisely, combining results from [11] and [32, 37], we establish a connection between the WL dimension and the alternating enumerator of certain graphs with respect to $\Phi$. We start by covering the necessary tools from [11] and [32, 37].
Definition 7.1.
A graph invariant $\Psi$ is a graph motif parameter if there is a finite collection of graphs ${\mathcal{L}}$ and coefficients $\alpha_{F}\in{\mathbb{R}}$ for all $F\in{\mathcal{L}}$ such that
$\Psi(G)=\sum_{F\in{\mathcal{L}}}\alpha_{F}\cdot\#\mathrm{Hom}(F\to G)$
for all graphs $G$.
The WL dimension of graph motif parameters has been determined in [32, 37].
Theorem 7.2 ([32, 37]).
Let $\Psi$ be a graph motif parameter and suppose there are a finite collection of pairwise non-isomorphic graphs ${\mathcal{L}}$ and coefficients $\alpha_{F}\in{\mathbb{R}}\setminus\{0\}$ for all $F\in{\mathcal{L}}$ such that
$\Psi(G)=\sum_{F\in{\mathcal{L}}}\alpha_{F}\cdot\#\mathrm{Hom}(F\to G)$
for all graphs $G$. Then the WL dimension of $\Psi$ is equal to $\max_{F\in{\mathcal{L}}}\operatorname{tw}(F)$.
Also, it has been shown in [11] that, for every graph invariant $\Phi$ on $k$-vertex graphs, the function #$\#\mathrm{IndSub}(\Phi\to\,\star\,)$ is a graph motif parameter. In fact, [11] even provides explicit formulas to determine the set ${\mathcal{L}}$ and the coefficients $\alpha_{F}$ for all graphs $F\in{\mathcal{L}}$. For the purpose of this paper, the following weaker version is sufficient.
For a graph $H$ consider a partition $\rho$ of $V(H)$ such that each block $B\in\rho$ is an independent set. Then we define $H/\rho$ as the graph obtained from $H$ by contracting each block to a single vertex. Formally, we set $V(H/\rho)=\rho$ and $E(H/\rho)=\{B_{1}B_{2}\mid\exists v_{1}\in B_{1},v_{2}\in B_{2}\colon v_{1}v_{2}\in E(H)\}$. We say a graph $F$ is a quotient of $H$, denoted by $F\preceq H$, if $F$ is isomorphic to $H/\rho$ for some partition $\rho$ of $V(H)$ such that each block $B\in\rho$ is an independent set.
Lemma 7.3 ([11]).
Let $\Phi$ be a $k$-vertex graph invariant. Then there is a finite collection of pairwise non-isomorphic graphs ${\mathcal{L}}$ and coefficients $\alpha_{F}\in{\mathbb{R}}\setminus\{0\}$ such that
  1. (I)
    #$\#\mathrm{IndSub}(\Phi\to G)=\sum_{F\in{\mathcal{L}}}\alpha_{F}\cdot\hom(F,G)$ for all graphs $G$,
  2. (II)
    $F\in{\mathcal{L}}$ for all graphs such that $\widehat{\Phi}(F)\neq 0$, and
  3. (III)
    if $F\in{\mathcal{L}}$, then there is some $k$-vertex graph $H$ such that $F\preceq H$ and $\widehat{\Phi}(H)\neq 0$.
The next lemma is an immediate consequence of Theorem 7.2 and Lemma 7.3.
Lemma 7.4.
Let $\Phi$ be a $k$-vertex graph invariant. Suppose there is a $k$-vertex graph $H$ with $\widehat{\Phi}(H)\neq 0$ and $\operatorname{tw}(H)\geq\ell$. Then the WL dimension of #$\#\mathrm{IndSub}(\Phi\to\,\star\,)$ is at least $\ell$.
We can combine Lemma 7.4 with Theorems 5.1, 5.5, 5.15, 5.20, 6.1 and 6.7 as well as Lemma 5.18 and Corollary 6.6 to obtain bounds on the WL dimension of certain graph invariants. Let us only explicitly mention the following consequence, which follows from Theorem 5.5 by setting $p=2$ and gives a lower bound that differs from the trivial upper bound of $k-1$ only by the constant factor $4$.
Corollary 7.5.
Let $k\geq 1$ and suppose $\Phi$ is a monotone $k$-vertex graph property. Then the WL dimension of #$\#\mathrm{IndSub}(\Phi\to\,\star\,)$ is strictly greater than $\frac{k}{4}$.
For $\Phi_{\sf even}$, we can even determine the WL dimension of #$\#\mathrm{IndSub}(\Phi_{\sf even}^{(k)}\to\,\star\,)$ exactly.
Theorem 7.6.
For every $k\geq 2$ the WL dimension of #$\#\mathrm{IndSub}(\Phi_{\sf even}^{(k)}\to\,\star\,)$ is equal to $\left\lfloor\frac{k}{2}\right\rfloor$.
The proof relies on the following simple lemma.
Lemma 7.7.
Let $H$ be a graph and $F$ a quotient of $H$. Then $\operatorname{tw}(F)\leq\operatorname{vc}(F)\leq\operatorname{vc}(H)$.
Proof.
The first inequality is folklore. For the second inequality, let $\rho$ be a partition of $H$ into independent sets, and let $C$ be a vertex cover of $H$. Then $C/\rho\coloneqq\{B\in\rho\mid B\cap C\neq\emptyset\}$ is a vertex cover of $H/\rho$. So $\operatorname{vc}(H/\rho)\leq\operatorname{vc}(H)$. Since $F$ is a quotient of $H$, the second inequality holds. ∎
Proof of Theorem 7.6.
Let $B\coloneqq K_{\left\lfloor k/2\right\rfloor,\left\lceil k/2\right\rceil}$. We have $\operatorname{tw}(B)=\left\lfloor\frac{k}{2}\right\rfloor$ and by Theorem 6.1, we have $\widehat{\Phi}_{\sf even}(B)\neq 0$. So the WL dimension of #$\#\mathrm{IndSub}(\Phi_{\sf even}^{(k)}\to\,\star\,)$ is at least $\left\lfloor\frac{k}{2}\right\rfloor$ by Lemma 7.4.
For the other direction, let $H$ be a graph such that $\widehat{\Phi}_{\sf even}(H)\neq 0$ and let $F$ be a quotient of $H$. Then $H$ is bipartite by Theorem 6.1 which implies that $\operatorname{tw}(F)\leq\operatorname{vc}(H)\leq\left\lfloor\frac{k}{2}\right\rfloor$ by Lemma 7.7. This implies that the WL dimension of #$\#\mathrm{IndSub}(\Phi_{\sf even}^{(k)}\to\,\star\,)$ is at most $\left\lfloor\frac{k}{2}\right\rfloor$ by Theorem 7.2 and Lemma 7.3(III). ∎
8 Conclusion
We show that relatively straightforward applications of Fourier analysis can yield strong complexity results for the problem of counting induced subgraphs satisfying a fixed property $\Phi$. Our mostly self-contained paper supersedes significant results for this problem obtained over the last decade via different techniques. Nevertheless, several open problems remain.
  1. 1.
    Most importantly, given a general non-meager property $\Phi$, is the parameterized problem #$\#\textnormal{{ParIndSub}}(\Phi)$ indeed #W[1]-hard? Note that this is not the case for all non-meager graph invariants $\Psi$. Indeed, consider the $k$-vertex graph invariant $\Psi$ that counts the copies of $K_{1,k-1}$ in a $k$-vertex graph. Then the problem #$\#\textnormal{{IndSub}}(\Psi)$ counts copies of $K_{1,k-1}$ in a graph $G$, which can be solved in linear time.
  2. 2.
    One approach towards settling the above question may be to invoke Corollary 6.9 for products $\Phi\cdot\Psi$ with other properties $\Psi$ that have at most moderately large degree, besides $\Phi_{\sf even}$. Given a $k$-vertex graph property $\Phi$, let $q_{k}$ for $k\in{\mathbb{N}}$ denote the maximum value of $\operatorname{tw}(H)$ among all $k$-vertex graphs $H$ with $\widehat{\Phi}(H)\neq 0$. Which values $q_{k}$ can be attained by a $k$-vertex graph property $\Phi$?
  3. 3.
    Finally, it is interesting to obtain tighter lower bounds for the Weisfeiler-Leman dimension of #$\#\mathrm{IndSub}(\Phi\to\,\star\,)$ for $k$-vertex graph invariants $\Phi$. Such bounds translate into tighter inexpressibility results in logics and computational lower bounds for symmetric circuits.
Appendix A Omitted Proofs
Graph invariants $f(\star)$ can be studied by viewing them as linear combinations of homomorphism counts $f(\star)=\sum_{i=1}^{s}\alpha_{i}\cdot\#\mathrm{Hom}(H_{i}\to\,\star\,)$ and exploiting a so-called complexity monotonicity property: In many important computational models, linear combinations of homomorphism counts as above are as hard to evaluate as their “hardest” homomorphism counts #Hom($\#\mathrm{Hom}(H_{i}\to\,\star\,)$ for $i\in[s]$, see [11].
In general, this does not hold for linear combinations $f(\star)=\sum_{i=1}^{s}\alpha_{i}\cdot\#\mathrm{Sub}(H_{i}\to\,\star\,)$ of subgraph counts. As an example, the easily computed quantity $\binom{E(G)}{k}$ counts all $k$-edge subgraphs of $G$. This count can be viewed as the sum of #$\#\mathrm{Sub}(H\to G)$ for all graphs $H$ with $|E(H)|=k$, which includes expanders $H$ on $\Omega(k)$ vertices and complete graphs on $\Omega(\sqrt{k})$ vertices. In this appendix, we show that a particular variant of the general complexity monotonicity property does however hold under the guarantee that all pattern graphs $H_{i}$ have the same number of vertices, and the pattern graph to be extracted is colorful: More precisely, if a graph $H$ appears with non-zero coefficient in the linear combination defining $f(\star)$, then there is an efficient reduction from counting colorful copies of $H$ in a vertex-colored graph $G$ to the evaluation of $f(\star)$ on uncolored graphs.
A.1 Vertex-Colored Graphs
We need to introduce some terminology for colored graphs that is not required outside of the appendix. In the following, a vertex-colored graph with colors $C$ is an undirected graph $G=(V,E,c)$ that is given together with a coloring $c\colon V(G)\to C$. Write $G^{\circ}=(V,E)$ for the uncolored graph underlying $G$. We call $G$ colorful if $c$ is a bijection, and we call it canonically colored if $C=V(G)$ and $c$ is the identity function.
For $i\in C$, we write $V_{i}(G)$ for the vertices of color $i$, and for $i,j\in C$, we write $E_{ij}(G)$ for the edges in $G$ with one endpoint of color $i$ and another of color $j$. For $X\subseteq C$ and $Y\subseteq\binom{C}{2}$, let $G_{\setminus X,Y}$ be the graph obtained from $G$ by deleting all vertices with colors from $X$ and all edges whose endpoints have a color pair from $Y$, i.e,
$G_{\setminus X,Y}=\left(V\setminus\bigcup_{i\in X}V_{i},\ E\setminus\bigcup_{ij\in Y}E_{ij}\right).$
Given vertex-colored graphs $H$ and $G$, we write #$\#\mathrm{Sub}(H\to G)$ for the number of subgraphs $F$ of $G$ that are isomorphic to $H$ by an isomorphism that respects colors. For colored graphs $F$ and $H$ with colorings $c_{F}$ and $c_{H}$, this is an isomorphism $\pi\colon V(F)\to V(H)$ with $c_{F}(v)=c_{H}(\pi(v))$ for all $v\in V(H)$. The following fact is immediate.
Fact A.1.
For graphs $H$ and $G$ with canonically colored $H$, we have
#$\#\mathrm{Sub}(H\to G)=\#\mathrm{Sub}(H\to G_{\setminus\emptyset,\overline{E(H)}}).$
Note that no vertices are removed, while all edges in $E_{ij}(G)$ with $ij\notin E(H)$ are removed. Thus, when considering #$\#\mathrm{Sub}(H\to G)$ for graphs $H$ and $G$ with canonically colored $H$, we may assume $E_{ij}(G)=\emptyset$ for all $ij\notin E(H)$.
A.2 Isolating Colorful Subgraphs from Linear Combinations
We follow the standard approach [10, 2] of applying the inclusion-exclusion principle on the set of subgraphs of a graph $G$. This principle allows us to express the number of subgraphs that are not contained in any bad sets $B_{1},\ldots,B_{s}$ as a signed sum over the counts of subgraphs in the intersections of these bad sets.
Lemma A.2.
Let $\Omega$ be a finite set and $B_{1},\ldots,B_{s}\subseteq\Omega$. For $X\subseteq[s]$, write $B_{X}\coloneqq\bigcap_{i\in X}B_{i}$ with the convention $B_{\emptyset}=\Omega$. Then
$\left|\Omega\setminus\bigcup_{i=1}^{s}B_{i}\right|=\sum_{X\subseteq[s]}(-1)^{|X|}\left|B_{X}\right|.$
Our main lemma shows how to evaluate #$\#\mathrm{Sub}(H\to G)$ for vertex-colored graphs $H$ and $G$, with colorful $H$, given access to a linear combination of subgraph counts $f(\star)$ that contains #$\#\mathrm{Sub}(H\to\,\star\,)$ with non-zero coefficient: It suffices to evaluate $f(G^{\prime})$ on $2^{|V(H)+|E(H)|}$ subgraphs $G^{\prime}$ of $G$ and take a signed sum of the values via the inclusion-exclusion principle.
We write $G\cong H$ to denote that two (colored or uncolored) graphs $G,H$ are isomorphic.
Lemma A.3.
Let $k\in\mathbb{N}$ and let the following be given:
  • Numbers $\alpha_{1},\ldots,\alpha_{s}\in\mathbb{Q}$ and pairwise non-isomorphic uncolored graphs $H_{1},\ldots,H_{s}$ with $|V(H_{i})|=k$ for all $i\in[s]$, which define the graph invariant
    $f(\star)\coloneqq\sum_{i=1}^{s}\alpha_{i}\cdot\#\mathrm{Sub}(H_{i}\to\,\star\,),$
  • a canonically colored graph $H$ with $V(H)=[k]$ and $H^{\circ}\cong H_{b}$ for some $b\in[s]$, and
  • a colored graph $G$ with coloring $c:V(G)\to[k]$ satisfying $E_{ij}(G)=\emptyset$ for $ij\notin E(H)$, see Fact A.1.
Then we have
$\alpha_{b}\cdot\#\mathrm{Sub}(H\to G)=\sum_{\begin{subarray}{c}X\subseteq V(H)\\ Y\subseteq E(H)\end{subarray}}(-1)^{|X|+|Y|}f(G_{\setminus X,Y}^{\circ}).$
Proof.
Since $E_{ij}(G)=\emptyset$ for $ij\notin E(H)$, every colorful subgraph in $G$ is isomorphic to a subgraph of $H$. We aim at applying the inclusion-exclusion principle. Towards this, given an uncolored graph $Q\subseteq H$ with $|V(Q)|=k$, we write
$\Omega(Q)=\{F\subseteq G^{\circ}\mid F\cong Q\}.$
Our aim is to isolate those $F\in\Omega(Q)$ satisfying the following conditions:
  • $F$ should be vertex-colorful under the coloring $c$ of $G$. This is equivalent to requiring $V_{i}(F)\neq\emptyset$ for all $i\in[k]$, since $F\cong Q$ has exactly $k$ vertices. For $i\in[k]$, we define the bad set $B_{i}(Q):=\{F\in\Omega(Q)\mid V_{i}(F)=\emptyset\}$ of subgraphs $F$ violating the condition for $i$.
  • $F$ should have $E_{ij}(F)\neq\emptyset$ for all $ij\in E(H)$. For $ij\in E(H)$, we define the bad set $B_{ij}(Q):=\{F\in\Omega(Q)\mid E_{ij}(F)=\emptyset\}$ violating the condition for $ij$.
Write $A(Q)$ for the subset of $\Omega(Q)$ that violates none of the conditions, that is,
$A(Q)\coloneqq\Omega(Q)\setminus\left(\bigcup_{i\in[k]}B_{i}(Q)\cup\bigcup_{ij\in E(H)}B_{ij}(Q)\right).$ (15)
Given a subgraph $F$ of $G^{\circ}$, we write $F^{c}$ for the colored graph obtained by applying $c$ to $F$. Every graph $F^{c}$ for $F\in A$ is colorful. Moreover, since $F$ is a subgraph of $G$ with $E_{ij}(G)=\emptyset$ for $ij\notin E(H)$, it is also isomorphic to a subgraph of $H$, and by the exclusion of graphs avoiding any $E_{ij}(G)$ for $ij\in E(H)$ from $A(Q)$, it is also isomorphic to a supergraph of $H$. Thus $F^{c}$ is isomorphic to $H$. This implies in particular that $F\cong Q\cong H^{\circ}$. We therefore have
$\{F^{c}\mid F\in A(Q)\}=\begin{cases}\{F\subseteq G\mid F\cong H\}&\text{if }Q\cong H^{\circ}\\ \emptyset&\text{otherwise}.\end{cases}$
Counting elements in the sets on both sides, it follows that
$|A(Q)|=\begin{cases}\#\mathrm{Sub}(H\to G)&\text{if }Q\cong H^{\circ}\\ 0&\text{otherwise}.\end{cases}$ (16)
Now we apply the inclusion-exclusion principle. We view $A(Q)$ as the set difference from (15) and express its cardinality via those of $B_{X,Y}\coloneqq\bigcap_{i\in X}B_{i}(Q)\cap\bigcap_{ij\in Y}B_{ij}(Q)$. Observe that $|B_{X,Y}|=\#\mathrm{Sub}(Q\to G_{\setminus X,Y}^{\circ})$. Invoking Lemma A.2, we obtain that
$\left|A(Q)\right|$ $=\sum_{\begin{subarray}{c}X\subseteq[k]\\ Y\subseteq E(H)\end{subarray}}(-1)^{|X|+|Y|}|B_{X,Y}|=\sum_{\begin{subarray}{c}X\subseteq[k]\\ Y\subseteq E(H)\end{subarray}}(-1)^{|X|+|Y|}\#\mathrm{Sub}(Q\to G_{\setminus X,Y}^{\circ}).$
Finally, we take a linear combination of the above equations for $Q\in\{H_{1},\ldots,H_{s}\}$, with coefficients $\alpha_{1},\ldots,\alpha_{s}$, and obtain
$\sum_{i=1}^{s}\alpha_{i}\left|A(H_{i})\right|=\sum_{i=1}^{s}\alpha_{i}\sum_{\begin{subarray}{c}X\subseteq[k]\\ Y\subseteq E(H)\end{subarray}}(-1)^{|X|+|Y|}\#\mathrm{Sub}(H_{i}\to G_{\setminus X,Y}^{\circ}).$
By (16), the left-hand side equals $\sum_{i=1}^{s}\alpha_{i}\left|A(H_{i})\right|=\alpha_{b}\cdot\#\mathrm{Sub}(H\to G)$. By the definition of $f(\star)$ and after changing the order of summation, we also observe that the right-hand side equals $\sum_{X\subseteq[k],\,Y\subseteq E(H)}(-1)^{|X|+|Y|}\cdot f(G_{\setminus X,Y}^{\circ})$. This proves the lemma. ∎
Before applying this lemma to obtain the complexity results in Theorems 3.4 and 3.5, we first derive Lemma 3.3 on the alternating enumerator from it.
Proof of Lemma 3.3.
Let $f(\star)=\sum_{H}\Phi(H)\cdot\#\mathrm{IndSub}(H\to\,\star\,)$. Write $F_{1},\ldots,F_{s}$ for $s\in{\mathbb{N}}$ for an enumeration of all graphs on vertex set $[k]$, with one graph per isomorphism class. It follows by linearity from [33, (5.17)] that there are coefficients $\alpha_{1},\ldots,\alpha_{s}\in{\mathbb{Q}}$ such that [^5] In the notation of [33], the expressions $\mathrm{ind}(H,G)$ and $\mathrm{inj}(H,G)$ are used. Translating into our notation, we have $\mathrm{ind}(H,G)=\#\mathrm{Aut}(H)\cdot\#\mathrm{IndSub}(H\to G)$ and $\mathrm{inj}(H,G)=\#\mathrm{Aut}(H)\cdot\#\mathrm{Sub}(H\to G)$.
$f(\star)=\sum_{i=1}^{s}\alpha_{i}\cdot\#\mathrm{Sub}(F_{i}\to\,\star\,).$
Given a graph $H$, let $H^{\textnormal{{can}}}$ be the canonically colored graph obtained from $H$. Let $b\in[s]$ be such that $H\cong F_{b}$. We have, using Lemma A.3 in the first equation,
$\alpha_{b}\cdot\underbrace{\#\mathrm{Sub}(H^{\textnormal{{can}}}\to H^{\textnormal{{can}}})}_{=1}$ $=\sum_{\begin{subarray}{c}X\subseteq V(H)\\ Y\subseteq E(H)\end{subarray}}(-1)^{|X|+|Y|}f(H_{\setminus X,Y})$
$=\sum_{Y\subseteq E(H)}(-1)^{|Y|}f(H_{\setminus\emptyset,Y})$
$=(-1)^{|E(H)|}\sum_{Y\subseteq E(H)}(-1)^{|Y|}f(H[Y]).$
In the second equation, we used that $f(F)=0$ if $|V(F)|<k$. In the third equation, we used that $H_{\setminus\emptyset,Y}=H[E(H)\setminus Y]$ and performed the index shift $Y\mapsto E(H)\setminus Y$ in the summation. This proves the lemma, as it shows that $\alpha_{b}$ indeed equals the alternating enumerator. ∎
A.3 Complexity Lower Bounds by Isolating Subgraphs
We are ready to use Lemma A.3 to obtain complexity lower bounds. First, the following useful fact allows us to restrict alternating enumerators to slices.
Fact A.4.
For every graph invariant $\Phi$ and $k\geq 1$, it holds that $\widehat{\Phi^{(k)}}(H)=\widehat{\Phi}(H)$ for all $k$-vertex graphs $H$.
Proof.
This follows directly from Definition 3.1 since $\widehat{\Phi}(H)$ depends only on the values of $\Phi$ on $k$-vertex graphs. ∎
We base our hardness results on the following parameterized problem: For a fixed class of uncolored graphs $\mathcal{H}$, the problem #$\#\textnormal{{ParColSub}}(\mathcal{H})$ asks, given as input $H\in\mathcal{H}$ and a vertex-colored graph $G$, to determine the number #$\#\mathrm{Sub}(H^{\textnormal{{can}}}\to G)$ where $H^{\textnormal{{can}}}$ denotes the canonically colored version of $H$. The problem is parameterized by $|V(H)|$.
Proof of Theorem 3.4.
[…]
Towards this end, let $\mathcal{H}=(H_{t})_{t\geq 1}$ be a graph sequence of unbounded treewidth and let $\Phi$ be a graph invariant with $\widehat{\Phi}(H_{t})\neq 0$ for all $t\geq 1$. We reduce #$\#\textnormal{{ParColSub}}(\mathcal{H})$ to #$\#\textnormal{{ParIndSub}}(\Phi)$ via Lemma A.3. Consider an instance $(H,G)$ for #$\#\textnormal{{ParColSub}}(\mathcal{H})$ and let $k\coloneqq|V(H)|$ and $n\coloneqq|V(G)|$. Using Fact A.1, we may assume that $G=G_{\setminus\emptyset,\overline{E(H)}}$. We wish to determine #$\#\mathrm{Sub}(H^{\textnormal{{can}}}\to G)$. By Lemma 3.3 we have
#$\#\mathrm{IndSub}(\Phi^{(k)}\to\star)=\sum_{F}\widehat{\Phi^{(k)}}(F)\cdot\#\mathrm{Sub}(F\to\star)=:f(\star),$
where $F$ ranges over all $k$-vertex graphs. Invoking Lemma A.3, we obtain that
$\widehat{\Phi^{(k)}}(H)\cdot\#\mathrm{Sub}(H\to G)=\sum_{\begin{subarray}{c}X\subseteq V(H)\\ Y\subseteq E(H)\end{subarray}}(-1)^{|X|+|Y|}f(G_{\setminus X,Y}^{\circ}).$ (17)
Since $\widehat{\Phi^{(k)}}(H)=\widehat{\Phi}(H)$ by Fact A.4 and $\widehat{\Phi}(H)\neq 0$ by assumption, we obtain a parameterized Turing reduction from #$\#\textnormal{{ParColSub}}(\mathcal{H})$ to #$\#\textnormal{{ParIndSub}}(\Phi)$ by evaluating the right-hand side of (17) and dividing by $\widehat{\Phi}(H)\neq 0$. Note that all relevant values $f(G_{\setminus X,Y}^{\circ})$ can be obtained by the oracle calls #$\#\mathrm{IndSub}(\Phi^{(k)}\to G^{\circ}_{\setminus X,Y})$ without parameter increase in overall time $2^{|V(H)|+|E(H)|}\cdot n^{O(1)}$. The value $\widehat{\Phi}(H)$ can be computed by brute-force by evaluating $\Phi$ on $2^{O(k^{2})}$ many $k$-vertex graphs. This yields the first statement of the theorem.
For the second statement, note that the above approach also yields a parameterized reduction from #$\#_{p}\textnormal{{ParColSub}}(\mathcal{H})$ to #$\#_{p}\textnormal{{ParIndSub}}(\Phi)$. In this setting, we evaluate the values $f(G_{\setminus X,Y}^{\circ})=\#\mathrm{IndSub}(\Phi^{(k)}\to G_{\setminus X,Y}^{\circ})$ modulo $p$ by oracle calls, evaluate the right-hand side of (17) modulo $p$, and then divide by $\widehat{\Phi}(H)$. Observe that $\widehat{\Phi}(H)\neq 0\bmod p$ by assumption, so the field ${\mathbb{Z}}_{p}$ with prime $p$ supports the division by $\widehat{\Phi}(H)$. ∎
[…]
Theorem A.5.
[…]
A known parameterized reduction from decision to modular counting [6, 49] allows us to derive the following consequence.
Theorem A.6.
[…]
Proof.
[…]
For the proof of Theorem 3.5, we also require the following lemma which is proved as a substatement in [10, Lemma 5.8] (see [10, (5.4)]).
Lemma A.7.
Let $F$ and $H$ be two graphs such that $F$ is a minor of $H$. Then there is a polynomial-time algorithm that, given a colored graph $G$, computes a colored graph $G^{\prime}$ such that
#$\#\mathrm{Sub}(F^{\textnormal{{can}}}\to G)=\#\mathrm{Sub}(H^{\textnormal{{can}}}\to G^{\prime}).$
We also rely on the following tool.
Theorem A.8 ([31]).
There is some constant $\delta>0$ such that every $n$-vertex graph with $|E(G)|\geq n\cdot\ell$ has a minor $K_{h}$ where
$h\coloneqq\left\lceil\delta\cdot\frac{\ell}{\sqrt{\log\ell}}\right\rceil.$
Proof of Theorem 3.5.
[…]
#$\#\mathrm{Sub}(K_{\ell}^{\textnormal{{can}}}\to G^{\textsf{col}})=\#\mathrm{Sub}(H^{\textnormal{{can}}}\to G^{\prime}).$
[…]
For Item (a) and the first part of Item (c), we proceed similarly. Let
$h\coloneqq\left\lceil\delta\cdot\frac{\ell}{\sqrt{\log\ell}}\right\rceil$
[…]
Next, we show an algorithmic upper bound omitted from the main part.
Proof of Theorem 3.6.
Let $(G,k)$ be an input for #$\#\textnormal{{ParIndSub}}(\Phi)$, where $G$ is a graph and $k\geq 1$ is an integer. We wish to compute #$\#\mathrm{IndSub}(\Phi^{(k)}\to G)$ for the $k$-th slice $\Phi^{(k)}$ of $\Phi$. Lemma 3.3 yields, with $H$ ranging over all unlabelled $k$-vertex graphs,
#$\#\mathrm{IndSub}(\Phi^{(k)}\to G)=\sum_{H}\widehat{\Phi^{(k)}}(H)\cdot\#\mathrm{Sub}(H\to G).$ (18)
Every graph $H$ with non-zero coefficient $\widehat{\Phi^{(k)}}(H)=\widehat{\Phi}(H)$ (recall Fact A.4 for the equality) satisfies $\operatorname{vc}(H)\leq t(k)$. Using Definition 3.1, all coefficients for all $k$-vertex graphs $H$ can be computed in time $2^{O(k^{2})}$ by brute-force. Thus, it remains to compute #$\#\mathrm{Sub}(H\to G)$ for each of the at most $2^{O(k^{2})}$ graphs $H$ with non-zero coefficient $\widehat{\Phi}(H)$; the non-zeroness of the coefficient implies $\operatorname{vc}(H)\leq t(k)$.
By [11, Theorem 1.1], the value #$\#\mathrm{Sub}(H\to G)$ on input graphs $H$ and $G$, with $k$ and $n$ vertices respectively, can be computed in time $f(k)\cdot n^{q(H)+1}$, where $q(H)\coloneqq\max_{F\preceq H}\operatorname{tw}(F)$ and $f$ is computable. Here $F\preceq H$ denotes that $F$ is a quotient of $H$, i.e., it can be obtained repeatedly identifying vertices; see also Section 7. Given a graph $H$ with $\operatorname{vc}(H)\leq t(k)$, we have $q(H)\leq t(k)$ by Lemma 7.7. Using [11, Theorem 1.1], it follows that an individual value #$\#\mathrm{Sub}(H\to G)$ for $H$ with $\widehat{\Phi^{(k)}}(H)=\widehat{\Phi}(H)\neq 0$ can be computed in time $f(k)\cdot n^{t(k)+1}$. Hence, #$\#\mathrm{IndSub}(\Phi^{(k)}\to G)$ can be computed in time $2^{O(k^{2})}f(k)\cdot n^{t(k)+1}$ via (18). ∎